### v1

# Score-based Generative Priors Guided Model-driven Network for MRI Reconstruction

Abstract
Score matching with Langevin dynamics (SMLD) method has been successfully applied to accelerated MRI. However, the hyperparameters in the sampling process require subtle tuning, otherwise the results can be severely corrupted by hallucination artifacts, particularly with out-of-distribution test data. In this study, we propose a novel workflow in which SMLD results are regarded as additional priors to guide model-driven network training. First, we adopted a pretrained score network to obtain samples as preliminary guidance images (PGI) without the need for network retraining, parameter tuning and in-distribution test data. Although PGIs are corrupted by hallucination artifacts, we believe that they can provide extra information through effective denoising steps to facilitate reconstruction. Therefore, we designed a denoising module (DM) in the second step to improve the quality of PGIs. The features are extracted from the components of Langevin dynamics and the same score network with fine-tuning; hence, we can directly learn the artifact patterns. Third, we designed a model-driven network whose training is guided by denoised PGIs (DGIs). DGIs are densely connected with intermediate reconstructions in each cascade to enrich the features and are periodically updated to provide more accurate guidance. Our experiments on different sequences revealed that despite the low average quality of PGIs, the proposed workflow can effectively extract valuable information to guide the network training, even with severely reduced training data and sampling steps. Our method outperforms other cutting-edge techniques by effectively mitigating hallucination artifacts, yielding robust and high-quality reconstruction results.
Index Terms:
MRI reconstruction, model-driven, diffusion model, compressed sensing, denoising score matching.
I Introduction
Magnetic Resonance Imaging (MRI) inherently requires long scanning times, leading to patient discomfort and degradation of image quality. Undersampling the k-space data can proportionally reduce the scanning time but results in aliasing artifacts in the reconstructed images. The conventional methods for reconstructing high-quality images from a small subset of k-space data are mainly parallel imaging (PI) [1, 2] and compressed sensing (CS) [3, 4, 5]. In recent years, deep-learning (DL)-based methods have been successfully introduced in the field of MRI reconstruction and have shown convincing improvements over conventional methods [6, 7]. For DL-based reconstruction, the proposed reconstruction models can be categorized into data-driven and model-driven unrolled networks. Part of the data-driven model was designed in an end-to-end manner to fit the mapping between the input measurements and reconstruction targets [8, 9, 10]. However, these methods still present problems such as lack of interpretability, and a relatively large dataset is required to fit the mapping. Recently, some data-driven networks have focused on distribution-learning-based reconstruction processes using diffusion models [11, 12, 13, 14, 15, 16, 17]. Among them, denoising diffusion probabilistic models (DDPMs) [18, 19] and score matching [20, 21] with Langevin dynamics (SMLD) methods, which correspond to Varience Preserving (VP) Stochastic differential equations (SDEs) and Variaence Exploding (VE) SDEs, respectively [22], have shown great potential in solving inverse problems. They have also been successfully extended to the MRI reconstruction field [23, 24, 22, 25]. However, the training (i.e., forward) and sampling (i.e., reverse) processes require careful tuning, otherwise the sampling results can be corrupted by hallucination artifacts, particularly with out-of-distribution test data.
Figure 1: Distribution-shift reconstruction results of a pretrained score network on a knee MRI dataset. The score network was trained on a brain MRI dataset. Images in odd columns are objective reality and those in even columns are sampling results.
Model-driven unrolled networks are built in a different manner: conventional optimization algorithms are unrolled and realized in deep networks, and stacked cascades are used to mimic the iterative steps in the optimization procedure. Compared with data-driven methods, unrolled networks can be better interpreted as those in which conventional priors can be appropriately incorporated, and they have shown convincing results in recent years. Although stacking more cascades can improve the reconstruction performance, large-scale backbone networks can hardly be applied to unrolled networks, considering the limitations of GPU memory, and a trade-off must be established. However, data- and model-driven methods attempt to recover the same target images along different paths, and the complementarity between these methods should be considered. Moreover, we observed the convincing potential of SMLD methods for MRI reconstruction. Current score-matching networks require retraining when applied to different datasets and we observe artifacts on the sampled images without network retraining and subtle tuning of hyperparameters, as shown in Fig. 1. However, we believe that the artifact-affected images can still provide information additional to the known undersampled measurements if the artifacts can be appropriately eliminated, thereby compensating for missing measurements.
Instead of focusing on optimizing the forward and reverse process in SMLD methods, we aim to design a model-driven network which can extract valuable information from naive sampling process and achieve high-quality results robustly. We proposed a novel DL-based hybrid MRI reconstruction workflow, consisting of the following three main steps: (1) Sampling: Using known undersampled raw k-space data, we first obtain samples using SMLD methods. We bypassed all the tuning steps in SMLD methods, such as tuning the hyperparameters or retraining the score networks on our dataset. Instead, we adopted a pretrained score network (PSN) and used the naive samples as preliminary guidance images (PGI) for the follow-up steps. A PSN can be trained using out-of-distribution data. (2) Denoising: We believe that these preliminary reconstructions can provide additional information to guide the reconstruction process through appropriate denoising, leading to a better performance. Considering that the PGIs have relatively low quality and that we observe similar aliasing artifacts in the images, we designed a denoising module (DM) to remove the artifacts from the PGIs. We extracted features from the same score network after fine-tuning and concatenated them with the inputs in Langevin dynamics, which directly produces artifacts, to learn the noise pattern. Then, we used denoised PGIs (DGIs) as higher-quality guidelines for further reconstruction. (3) Guided reconstruction: We designed a densely connected unrolled network to reconstruct the MR images from the undersampled measurements. DGIs are simultaneously input with measurements and periodically updated in the network to guide network training. The DM and unrolled cascades were trained end-to-end.
The experimental results show that although the quality of PGIs is even lower than that of the undersampled measurements, the artifacts were effectively eliminated by the DMs. Meanwhile, the score network can not only help in the sampling process but also cooperate with the unrolled network and improve the denoising performance of the DMs. The proposed reconstruction workflow outperformed state-of-the-art methods through the guidance of DGIs.
The main contributions of this study are summarized as follows:
  1. 1.
    We propose a novel DL-based MRI reconstruction workflow, and low-quality samples obtained using SMLD method are used as guidance during reconstruction.
  2. 2.
    We design an unrolled network, and the initially obtained samples are denoised and periodically updated to guide network training.
  3. 3.
    Experimental results show that without tuning the sampling steps or retraining the score network on target dataset, the low-quality samples can guide the proposed network to obtain robust and high-quality reconstructions, which outperforms the cutting-edge methods.
II Related Works
II-A Conventional approaches for accelerated MRI
For the PI, multiple receiver coils are equipped to obtain raw k-space data simultaneously, and the missing data are recovered by exploiting redundancies among the data from different coils. Sensitivity encoding (SENSE)-type [26] methods work in the image domain, whereas others directly interpolate missing data points in k-space, such as simultaneous acquisition of spatial harmonics (SMASH) [27] and generalized autocalibration partial parallel acquisition (GRAPPA) [28]. However, the reconstructions in PI approaches deteriorate at high acceleration factors. In CS, sparsity-based priors are routinely incorporated to narrow the solution space in the objective functions, and iterative optimization algorithms are designed to approach artifact-free images [29, 30, 31, 32, 33, 34]. However, conventional methods depend heavily on handcrafted procedures, and the iterative steps are time-consuming.
Figure 2: (a) The workflow of the proposed SGM-Net. (b) The network structure of the sampling and denoising steps. (c) The network structure of the $k$th cascade.
II-B Data-driven networks for MRI reconstruction
For data-driven networks trained end-to-end, the mapping between the undersampled measurements and reconstructed images was established. For example, early studies adopted deep convolutional neural networks (CNNs) to reconstruct MR images from zero-filled inputs [35, 36]. The authors in [37] trained the magnitude and phase networks separately in a deep residual network to remove aliasing artifacts. In [38], a complex-valued deep residual network with a higher performance and smaller size was proposed. In addition, GAN-[39, 40]-based networks have shown convincing results [41, 42, 43]. In addition to realizing image domain end-to-end learning, some studies have proposed direct learning mapping in k-space [44] or exploiting the complementarity between the image and k-space domains [9].
Distribution learning is another category of data-driven networks. For example, in [45] the authors used pre-trained generative models to present data distributions. In [23], the authors used score matching to estimate the data distribution, and samples were then generated using annealed Langevin dynamics. Subsequently, the authors proposed improved skills in [24] for higher stability and better performance, and exact log-likelihood computation was supported in [22]. [25] first successfully applied CS-based generative models to overcome the MRI reconstruction problem and showed convincing results in out-of-distribution samples. In [16], the authors further optimized the sampling stage by automatically tuning the hyperparameters using Stein’s unbiased risk estimator (SURE). The authors in [17] proposed a high-frenquency-based diffusion process and accelerated the reverse process. A wavelet-improved technique was proposed in [15] to stably train the score network. In the reverse process, the authors further designed regularization constraints to enhance the robustness. Some works extended the idea of cold diffusion [46] to MRI reconstruction, and the process of adding Gaussian noise is replaced with undersampling operation [47]. However, network retraining is necessary when changing the acquisition parameters, and large datasets are required to fit the mapping [48].
II-C Unrolled networks for MRI reconstruction
Unrolled networks, also known as model-driven networks, are constructed by unrolling iterative optimization steps into deep networks [49, 50, 51, 52, 53]. For example, algorithms such as the alternating direction method of multipliers (ADMM) [49], iterative shrinkage-thresholding algorithm (ISTA) [50], variable splitting (VS) [52], and approximate message passing (AMP) have been successfully applied to solve the objective functions in MRI reconstruction. In [53], the authors extended a previous variational network [51] to learn end-to-end, and sensitivity maps (SM) were estimated in the network. To further improve the accuracy of the SM, different studies [54, 55] have explored ways to simultaneously update the SMs with the reconstructed images. In [56], convolutional blocks were used to extract high-resolution features and transformer blocks have been introduced to refine low-resolution features. The authors of [57] used pre-acquired intra-subject MRI modalities to guide the reconstruction process. In [58], the authors designed a transformer-enhanced network, incorporating a regularization term on the error maps within the residual image domain. Some works solved the optimization problem with structured low-rank algorithm [59, 60, 61]. The authors in [62] designed a deep low-rank and sparse network which explored 1D convolution during training. However, the complementarity between unrolled networks and data-driven methods has not been investigated; in this study, we propose a novel concept in which we treat naive sampling results from SMLD methods as guidance and unrolled cascades as denoising and refining processes.
III Background
In PI, the acquisition of multicoil measurements can be formulated as
$y=\mathcal{A}x+b$ (1)
where $y\in\mathbb{C}^{NC}$ is the undersampled k-space data and $C$ is the number of coils. $x\in\mathbb{C}^{M}$ is the underlying MR image for reconstruction and $N<M$. $b$ is the measurement noise. $\mathcal{A}$ is the forward operator given by
$\mathcal{A}(\cdot)=\mathcal{P}\circ\mathcal{F}\circ\mathcal{E}(\cdot)$ (2)
where $\mathcal{P}$ denotes the sampling masks that zero the undersampled data points, and $\mathcal{F}$ denotes the Fourier transform matrix. The expanding operator $\mathcal{E}(\cdot)$ calculates the coil-specific images from $x$, defined as
$\mathcal{E}(x)=(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...),i=1,2,...,C$ (3)
where $\mathcal{S}_{i}$ is the SM of the $i$th coil. Conversely, the reduced operator $\mathcal{R}$ integrates the coil-specific images into a single image defined as follows:
$\mathcal{R}(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...)=\sum_{i=1}^{C}\mathcal{S}_{i}x$ (4)
and the backward operator is expressed as
$\mathcal{A}^{*}(\cdot)=\mathcal{R}\circ\mathcal{F}^{-1}\circ\mathcal{P}(\cdot)$ (5)
It is ill-posed to recover missing data directly from $y$. Hence, a regularization term $\lambda\mathcal{U}(x)$ is routinely added to narrow the solution space in the CS, which is
$\hat{x}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{C}\|\mathcal{P}\mathcal{F}\mathcal{S}_{i}x-y\|_{2}^{2}+\lambda\mathcal{U}(x)$ (6)
For example, $\lambda\Psi(x)$ is a common regularization term that imposes the sparsity constraint. Then, gradient descent can be used to solve the problem iteratively as follows:
$x^{t+1}=x^{t}-\alpha^{t}(\mathcal{A}^{*}(\mathcal{A}x^{t}-y)+\lambda\psi(x^{t}))$ (7)
where $\alpha^{t}$ is the step size in the $t$th iteration and $\psi(x^{t})$ is the gradient of $\Psi(x)$.
IV Methods
We proposed a score-based generative priors guided model-driven Network (SGM-Net), which is illustrated in Fig. 2. The whole workflow is illustrated in Fig. 2 (a), which comprises three parts: sampling (green components), denoising (blue components), and guided reconstruction (yellow components). The preliminary samples (that is, $x_{T}$ in Fig. 2 (a)) were first obtained using Langevin dynamics and then used as PGIs for the follow-up steps. The gradients of the data distribution were obtained using PSN. Instead of carefully training and tuning the SMLD process to obtain high-quality samples, we sampled $x_{T}$ from an out-of-distribution pretrained network with fitted hyperparameters. Then, we treat the artifacts in $x_{T}$ as removable and propose a DM to eliminate noise and enhance $x_{T}$. The information available from the sampling steps was integrated and fed into the DM, as shown in Fig. 2 (b). Finally, DGIs were used to guide an unrolled network, as shown in Fig. 2 (c). Details of the proposed method are presented in the following subsections.
IV-A Score matching and Langevin dynamics
The first step of the proposed workflow is to obtain PGIs with SMLD to guide the subsequent steps. In this study, we adopted the PSN proposed in [23], which was successfully extended to MRI reconstruction in [25]. To improve the generality of our workflow, the score network was pretrained with out-of-distribution data and tedious tuning and retraining steps were bypassed. The network, remarked as $s_{\theta}$, is a four-cascade RefineNet [63] with dilated convolutions and instance normalization, which is used to approach the score function of $p(x)$ (that is, $\bigtriangledown_{x}\log p(x)$) and we have $s_{\theta}\approx\bigtriangledown_{x}\log p(x)$. Average pooling layers were adopted to replace all maximum pooling layers in RefineNet, and ELU was adopted as the activation function. Then, the annealed Langevin method-based sampling process in MRI can be represented as
$x_{t+1}=x_{t}+\eta_{t}(s_{\theta}(x_{t};\sigma_{t})+\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{t})}{\gamma_{t}^{2}+\sigma^{2}})+\sqrt{2\eta_{t}}\xi_{t}$ (8)
where $\xi_{t}\sim\mathcal{N}(0,\textit{I})$ and $t=0,1,2,..,T$. $\eta_{t}$ is the step size in the $t$th iteration and $\eta_{t}=\epsilon\cdot\frac{\sigma^{2}_{j}}{\sigma^{2}_{L}}$. $\{\sigma_{j}\}_{j=1}^{L}$ is a sequence of noise scales that gradually reduces each step size in the annealed Langevin dynamics, where $\sigma_{1}>\sigma_{2}>\cdots>\sigma_{L}$. $\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{t})}{\gamma_{t}^{2}+\sigma^{2}}$ is the consistency term introduced to exploit known measurements. Finally, $x_{T}$ approaches the target as $T$ becomes sufficiently large, which is used as a guide in the follow-up steps.
IV-B Denoising module
Over-smoothing details and hallucination artifacts were observed on the PGIs obtained in the first step, and the denoising step aimed to initially eliminate artifacts and compensate for missing details. First, we designed a KI-branch to learn an denoising mapping from the data between the frequency domain and the image domain. Considering that artifacts usually have different frequencies, which makes it possible to separate them in the k-space, we transferred $x_{T}$ back to the k-space and used a U-Net with a skip connection to learn the mapping to the denoised outputs. Subsequently, U-Net learned to eliminate the noisy signals in k-space. A data consistency (DC) block was concatenated to the U-Net to maintain consistency with the measurements. In addition, the complementarity between the k-space and image domain could also help obtain better performance. We then used an inverse Fourier transform to transfer the k-space data back to the image domain, which is denoted as $x_{d}$, as shown in Fig. 2 (b).
Then, we designed an input-enriched U-Net, marked as $N_{d}$, to further eliminate the noises in $x_{d}$. Fundamentally, the score network is used to generate clean images from noises and the artifacts are generated during the sampling process. Consequently, there could exist an underlying mapping between the artifacts and the elements involved in the sampling process (i.e. equation (8)). Therefore, with the input $x_{T}$ in the DM, we fed $x_{T}$, $s_{\theta}(x_{T};\sigma_{T}$) and $\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{T})}{\gamma_{T}^{2}+\sigma^{2}}$ as the input elements. We fine-tuned $s_{\theta}$, marked as $f_{s_{\theta}}$ that we removed the last batch normalization and activation layers to avoid undesirable loss of information, and concatenated a fine-tuning convolutional (FTC) layer to learn the noise pattern from the features. Only the FTC layer in $f_{s_{\theta}}$ was updated during training. The parameters in the other parts of $f_{s_{\theta}}$ were fitted, and the Langevin sampling results were fed to $f_{s_{\theta}}$ such that the FTC layers could directly learn the noise pattern from the score-matching process for each image.
Besides, to avoid newly introduced noises when obtaining $x_{d}$ from $x_{T}$, $x_{d}$ was also fed to $N_{d}$. Therefore, the enriched input features of $N_{d}$ have four components. By concatenating and feeding the four components into the DM, the denoising process is denoted as:
$x_{T}^{0}=x_{d}+N_{d}\{x_{p},x_{T},\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{T})}{\gamma_{T}^{2}+\sigma_{T}^{2}},x_{d}\}$ (9)
[…]
Figure 3: Examples of ground truth images (the first column), undersampled images (the second column) and the corresponding k-space trajectories (the third column). The images from the first and second rows are selected from the Knee dataset with four-fold and six-fold random Cartesian mask. The images from the third and fourth row are selected from the fastMRI dataset with four-fold and six-fold equispaced Cartesian mask.
coronal PD 4$\times$ acceleration
coronal PD 6$\times$ acceleration
FastMRI 4$\times$ acceleration
FastMRI 6$\times$ acceleration
Figure 4: Examples of the reconstructed images of different methods from coronal PD sequence and FastMRI dataset at 4$\times$ and 6$\times$ acceleration. The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps corresponding to the objective reality. The methods from left to right: 1. objective reality; 2. zero-filled; 3. TV; 4. U-Net; 5. D5C5; 6. ISTA-Net; 7. VS-Net; 8. E2EVarNet; 9. ReVarNet; 10. MeDL-Net; 11. SGM-Net (ours).
IV-C Unrolled network with PGI-guided densely-connected cascades
The third step of the proposed workflow is to design an unrolled network to obtain higher-quality reconstructions from denoised guidance images (DGIs) (that is, $x_{T}^{0}$), and the undersampled measurements $x_{z}^{0}$ and $x_{T}^{0}$ are used as two different constraints during unrolled network training. As shown in Fig. 2, the network consists of several DGI-guided densely connected cascades (GICs). The inputs of GICs are $x_{T}^{k}$ and $x_{z}^{k}$ where $k=0,1,...,K$, and $x_{z}^{0}$ is the coil-combined zero-filled image of the undersampled measurements. In each GIC, the undersampled measurements $x_{z}^{k}$ were periodically updated to approach the target images, whereas the DGIs $x_{T}^{k}$ were simultaneously updated to provide more accurate guidance, forming a parallel network structure. $I$ Data fidelity (DF) and $I$ U-Net blocks were alternately stacked in each GIC to solve (7), and U-Net blocks were used to learn the gradient of $\Psi(x)$ from the data. For the $x_{T}^{k}$ branch, the calculations in DF blocks can be represented as
$r_{T}^{k_{i+1}}=x_{T}^{k_{i}}-\alpha^{k_{i}}(\mathcal{A}^{*}(\mathcal{A}x_{T}^{{k_{i}}}-y)),i=0,1,...,I-1$ (10)
where $x_{T}^{k_{0}}$ is equal to $x_{T}^{k}$. For the $x_{z}^{k}$-branch, the DF blocks are given by
$r_{z}^{k_{i+1}}=x_{z}^{k_{i}}-\alpha^{k_{i}}(\mathcal{A}^{*}(\mathcal{A}x_{z}^{k_{i}}-y)),i=0,1,...,I-1$ (11)
where $x_{z}^{k_{0}}$ is equal to $x_{z}^{k}$. In addition, we propose a densely connected structure to enrich the input information for each U-Net block. For the $x_{T}^{k}$-branch, the initial input and each output of DF blocks corresponding to a U-Net block are concatenated before being fed into U-Net. Therefore, we have:
$x_{T}^{k_{i+1}}=U_{T}^{i}(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}}\})$ (12)
where $U_{T}^{i}$ denotes the $i$th U-Net block in the $x_{T}^{k}$-branch, $x_{T}^{k}$ denotes the input to the corresponding branch, and $r_{T}^{k_{i}}$ denotes the output of the $i$th DF block. A similar strategy is implemented in the $x_{z}^{k}$-branch. Considering that we take the final output of the $x_{z}^{k}$ branch in the last GIC as the final reconstruction, the $x_{z}^{k}$ branch is essentially the main branch, and the $x_{T}^{k}$ branch is simultaneously used to guide the reconstruction in the main branch. Therefore, the concatenated information from the $x_{T}^{k}$ branch is also fed to the $x_{z}^{k}$ branch to learn from the features from different stages, represented as
$x_{z}^{k_{i+1}}=U_{z}^{i}(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\})$ (13)
After the DF and U-Net blocks, we designed a skip connection and a DC block at the end of the $x_{T}^{k}$ branch, and the produced $x_{T}^{k+1}$ is fed to the next GIC as one of the two inputs as follows:
$x_{T}^{k+1}=DC(x_{T}^{k}+x_{T}^{k_{I}})$ (14)
For $x_{z}^{k}$, we further adopted a four-layer attention module (AM) [64] between the last U-Net and skip connection. The information fed to the last DF block is concatenated and simultaneously fed to the AM to produce an attention map
$M_{A}=AM(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\})$ (15)
Through a skip connection and DC block, $x_{z}^{k+1}$ is given as
$x_{z}^{k+1}=DC(x_{z}^{k}+x_{z}^{k_{I}}\odot M_{A}+x_{T}^{k_{I}}\odot((\textbf{J}-M_{A}))$ (16)
where J is an all-ones matrix and $\odot$ denotes the Hadamard product. Element-wise attention results in a deeper fusion between the two branches and optimizes the intermediate results in each $x_{z}^{k}$ branch.
IV-D loss functions
In this study, the loss function for unrolled network training was designed as
$Loss=\gamma_{1}(x_{d}+x_{T}^{0}+x_{z}^{K})+\gamma_{2}(\sum_{k=1}^{K}x_{T}^{k}+\sum_{k=1}^{K}x_{z}^{k})$ (17)
where $x_{d}$ amd $x_{T}^{0}$ are important guidelines for ensuring reconstruction performance, and $x_{z}^{K}$ is the result. $\sum_{k=1}^{K}x_{T}^{k}$ and $\sum_{k=1}^{K}x_{z}^{k}$ are intermediate results. We set $\gamma_{1}=1$ and $\gamma_{2}=0.1$ to distribute more weights for the guidance and results and fewer weights to help the network converge.
IV-E Training details
We adopted the score network $s_{\theta}$ released by the authors in [25], which is pretrained on a brain dataset. We followed the proposed hyperparameters in that manuscript [^1] https://github.com/utcsilab/csgm-mri-langevin during sampling, and we do not retrain the network on our knee dataset. The numbers of the four-cascade filters were 128, 256, 512, and 1024. The number of filters in the FTC layer is set to 2. In SGM-Net, $K$ is set to five and $I$ is set to three. The numbers of filters in all U-Nets in the DM and GICs were set to 16, 32, 64, and 128. The step size in DF blocks was learned directly in the network. All complex-valued MR images were stored in two channels during the network training. In particular, the first U-Net in DM uses k-space data as input, and the coil channel is combined with the feature channel. Therefore, the number of input and output channels of U-Net was set to 30 (15 coils). The proposed SGM-Net was trained on a UBUNTU 20.04 LTS system with a 3090 GPU (24GB) for 100 epoches, and the learning rate was set to 1e-3. The batch size is set to 1.
V Experiments and Results
V-A Dataset
TABLE I: The quantitative results on Coronal PD and PDFS sequences, as well as the FastMRI dataset, compared with those of cutting-edge methods.
Sequence Method PSNR SSIM
Coronal PD zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
TV 33.7498$\pm$3.1753 32.3290$\pm$3.2159 0.8851$\pm$0.0657 0.8658$\pm$0.0725
U-Net 36.8984$\pm$2.7072 34.7352$\pm$2.7863 0.9361$\pm$0.0610 0.9133$\pm$0.0682
D5C5 38.7903$\pm$2.8408 35.9697$\pm$2.7990 0.9509$\pm$0.0587 0.9259$\pm$0.0689
ISTA-Net 39.2910$\pm$2.8423 35.5220$\pm$3.0915 0.9471$\pm$0.0567 0.9138$\pm$0.0717
VS-Net 40.1649$\pm$2.8767 37.0604$\pm$2.7927 0.9591$\pm$0.0569 0.9358$\pm$0.0648
E2E-VarNet 39.3592$\pm$2.9368 36.8949$\pm$3.3872 0.9542$\pm$0.0662 0.9342$\pm$0.0775
ReVarNet 40.1141$\pm$2.8765 37.3681$\pm$2.7867 0.9590$\pm$0.0588 0.9379$\pm$0.0677
MeDL-Net 40.9295$\pm$2.9072 38.1052$\pm$2.7622 0.9614$\pm$0.0557 0.9417$\pm$0.0663
SGM-Net (ours) 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
Coronal PDFS zero-filled 32.0299$\pm$2.0273 31.2221$\pm$2.0575 0.7964$\pm$0.0998 0.7587$\pm$0.1172
TV 33.7256$\pm$1.9058 32.7328$\pm$1.8850 0.8124$\pm$0.0752 0.7920$\pm$0.0796
U-Net 34.4367$\pm$0.6572 33.2028$\pm$2.5662 0.8183$\pm$0.0106 0.7758$\pm$0.1222
D5C5 34.9671$\pm$2.8261 33.4189$\pm$2.4765 0.8250$\pm$0.1106 0.7854$\pm$0.1223
ISTA-Net 34.4990$\pm$2.6389 33.4569$\pm$2.5758 0.8198$\pm$0.1066 0.7854$\pm$0.1214
VS-Net 34.6847$\pm$2.7134 33.5010$\pm$2.6100 0.8259$\pm$0.1073 0.7842$\pm$0.1229
E2E-VarNet 34.6274$\pm$2.7331 33.1626$\pm$2.4742 0.8234$\pm$0.1101 0.7804$\pm$0.1244
ReVarNet 34.8522$\pm$2.7461 33.3146$\pm$2.5604 0.8272$\pm$0.1080 0.7832$\pm$0.1232
MeDL-Net 35.5931$\pm$2.7307 34.1409$\pm$2.5501 0.8349$\pm$0.1025 0.7922$\pm$0.1255
SGM-Net (ours) 35.7932$\pm$3.0718 34.3309$\pm$2.5192 0.8413$\pm$0.1076 0.7978$\pm$0.1242
FastMRI zero-filled 33.2828$\pm$3.0594 32.5069$\pm$3.0657 0.8857$\pm$0.0709 0.8680$\pm$0.0791
TV 35.5512$\pm$2.9189 33.8441$\pm$2.8703 0.8884$\pm$0.0635 0.0201$\pm$0.0152
U-Net 35.21158$\pm$2.5925 33.7864$\pm$2.5331 0.8562$\pm$0.0574 0.8091$\pm$0.0637
D5C5 36.4550$\pm$3.0246 34.5636$\pm$2.8510 0.9113$\pm$0.0649 0.8849$\pm$0.0749
ISTA-Net 36.5753$\pm$2.9887 34.6229$\pm$2.8269 0.9134$\pm$0.0656 0.8837$\pm$0.0752
VS-Net 37.3042$\pm$3.1130 35.5335$\pm$2.8513 0.9202$\pm$0.0637 0.8916$\pm$0.0736
E2E-VarNet 37.5951$\pm$3.0129 35.6696$\pm$2.8326 0.9217$\pm$0.0635 0.8978$\pm$0.0726
ReVarNet 37.0047$\pm$3.0130 34.8855$\pm$2.7926 0.9168$\pm$0.0646 0.8895$\pm$0.0746
MeDL-Net 37.6791$\pm$2.9644 35.9594$\pm$2.7401 0.8984$\pm$0.0643 0.8882$\pm$0.0717
SGM-Net (ours) 38.9212$\pm$3.4296 37.9430$\pm$3.2521 0.9344$\pm$0.0604 0.9234$\pm$0.0067
In this study, we adopted a publicly available 15-coil knee dataset proposed in [51] to train our model and compared it with other methods. The knee dataset consisted of five 2D turbo spin-echo (TSE) sequences, each of which contained volumes scanned using a 3T scanner from 20 patients. We conducted experiments on coronal proton density (PD; TR=2750 ms, TE=27 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 35-42 slices), and coronal fat-saturated PD (PDFS; TR=2870 ms, TE=33 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 33-44 slices) sequences. The coil sensitivity maps were precomputed using ESPRIRiT and included in the dataset. The training, validation, and testing sets were randomly divided at a ratio of 3:1:1. Random Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration. 8% lines at the center region are preserved. Examples of training images and the corresponding k-space trajectories are shown in Fig. 3 (the first and second rows).
FastMRI [8] is a publicly available dataset for accelerated MRI, which has been adopted in relevant studies. Considering the limitations of our hardware, we randomly selected a subset of the 15-coil knee dataset to evaluate our model and demonstrate its effectiveness. The data were acquired using the 2D TSE protocol, containing PD and PDFS sequence scanning from 3T or 1.5T systems (TR=2200-3000 ms, TE=27-34 ms, in-plane resolution 0.5$\times$0.5 mm2, slice thickness 3 mm). A total of 419, 33, and 375 slices were used for training, validation, and testing sets, respectively. Coil sensitivity maps were not available in the dataset, and related studies have been conducted to estimate and optimize sensitivity maps during training. However, we did not focus on SM estimation in this study and used pretrained SMs in our model to highlight the improvements brought about by the proposed methods. In the fastMRI dataset, we calculated the SMs using the sigpy [65] toolbox. Equispaced Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration. 8% lines at the center region are preserved. Examples of training images and the corresponding k-space trajectories are shown in Fig. 3 (the third and fourth rows).
V-B comparisons with the cutting-edge methods
We compared the proposed SGM-Net with two traditional and seven DL-based methods, as presented in Table I. In the zero-fill method, the missing data in the measurements were filled with zero. The TV method uses total variation as the image prior in a CS-based approach. For DL-based methods (i.e., U-Net [8] [^2] https://github.com/facebookresearch/fastMRI, D5C5 [66] [^3] https://github.com/js3611/Deep-MRI-Reconstruction. ISTA-Net [50] [^4] https://github.com/jianzhangcs/ISTA-Net-PyTorch VSNet [52] [^5] https://github.com/j-duan/VS-Net. E2EVarNet [53] [^6] https://github.com/NKI-AI/direct, ReVarNet [67]6 and MeDL-Net [68] [^7] https://github.com/joexy312/MEDL-Net), we adopted the released codes and retrained the networks on our device. We made a concerted effort to keep the announced settings same as those in the corresponding articles, except that we reduced the number of filters in ReVarNet so that all networks could be trained on one GPU. All methods were implemented on the knee dataset (coronal PD and PDFS sequences) and a subset of fastMRI dataset. The PSNR and SSIM results were calculated under four- and six-fold acceleration to evaluate the different methods. As shown in Table I, the proposed SGM-Net robustly achieves the best results on all the datasets with different acceleration factors. The reconstruction results of the sample images are presented in Fig.4. We can observe from the zoomed region (second rows) that our method recovered the most details, and the error maps (third row) exhibited the least errors.
V-C Ablation Studies
TABLE II: Results of ablation studies on coronal PD sequence with four- and six-fold acceleration rates.
Method PSNR SSIM
SG 26.8198$\pm$4.7495 25.0439$\pm$7.2973 0.4172$\pm$0.1712 0.5212$\pm$0.1568
MN 38.8362$\pm$2.6402 35.9488$\pm$2.6333 0.8623$\pm$0.0880 0.7861$\pm$0.1088
SR-MN 35.4585$\pm$2.9562 33.2093$\pm$2.6398 0.8351$\pm$0.0912 0.7412$\pm$0.1284
SG-MN-DM-US 40.7095$\pm$3.0165 38.0227$\pm$2.9647 0.9619$\pm$0.0615 0.9440$\pm$0.0700
SG-MN-DM-DG 40.8443$\pm$3.0384 38.2183$\pm$2.9986 0.9629$\pm$0.0624 0.9448$\pm$0.0721
SG-MN-US-DG 41.0424$\pm$3.4144 38.1616$\pm$2.9985 0.9633$\pm$0.0676 0.9443$\pm$0.0714
SG-MN-DM-US-DG 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
To highlight the improvements brought by different modules, we designed several workflows for a clear evaluation. Specifically, the proposed modules for evaluation were the SMLD-based guidance (SG), model-driven network structure (MN), denoising module (DM), updating steps (US) for SGs, and densely connected guidance (DG) in the unrolled network. The ablations was designed as follows.
  1. 1.
    SG: using the SMLD results as the final reconstructions.
  2. 2.
    MN: a standard unrolled network with the undersampled measurements as the inputs.
  3. 3.
    SG-MN: a standard unrolled network with SMLD results as the inputs.
  4. 4.
    SG-MN-DM-US: the densely connected structure is removed from all the cascades.
  5. 5.
    SG-MN-DM-DG: the SMLD results are used as guidance in the unrolled network but without and updating steps.
  6. 6.
    SG-MN-US-DG: the SMLD results are used to guide the network without denoising in the DM.
  7. 7.
    SG-MN-DM-US-DG: the proposed SGM-Net.
The quantitative results are shown in table II. Comparing SG with the proposed SGM-Net, we can observe a significant improvement brought about by the proposed modules, whereas the average results from the SMLD sampling are worse than those of the zero-filled method. Similarly, through the guidance of the SMLD results, the results of the unrolled MN method were significantly improved with guidance from the SMLD results. Using either of these two methods alone can’t achieve satisfactory results. Moreover, we simply combine SR and MN in the SR-MN workflow that the SMLD results are directly fed to the network as input. However, the performance are even worse than that of MN. This phenomenon demonstrates the effectiveness of the proposed guidance-based workflow, which outperforms a simple combination of the two methods.
A comparison between SG-MN-DM-US and SGM-Net demonstrated that the densely connected structure is capable of fulfilling the guiding work and improving the final results. Removing the densely connected structure and using only the AMs to optimize the reconstructed images will lead to a degredation in the final results. By comparing SG-MN-DM-DG with SGM-Net, we observed that the updating steps in the proposed model can ensured that accurate guidance was provided by the PGIs and improved the results. When removing the DM and directly feeding SMLD results to guide the network, the reconstruction performances of SG-MN-US-DG are similarly worse than that of SGM-Net. The best results are achieved with the integration of all the proposed modules, which demonstrates their effectiveness and necessity. Furthermore, we illustrate in subsection V.E that the integration of the proposed module can enhance the robustness of the network with greatly reduced training data.
Figure 5: The reconstruction results of the original model at four-fold acceleration. The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps representing the objective reality. The corresponding images from left to right: the objective reality, zero-filled measurements, PGI, DGI, and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branch.
Figure 6: The reconstruction results with 20% sampling steps at four-fold acceleration. The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps representing the objective reality. The corresponding images from left to right: the objective reality, zero-filled measurements, PGI, DGI, and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branch.
TABLE III: Quantitative results of the selected images at four-fold and six-fold acceleration rates.
Method Image PSNR SSIM
zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
original $x_{T}$ 26.8198$\pm$4.7495 25.0439$\pm$7.2973 0.4172$\pm$0.1712 0.5212$\pm$0.1568
$x_{T}^{0}$ 39.0871$\pm$2.5214 37.0431$\pm$2.3824 0.9523$\pm$0.0587 0.9364$\pm$0.0635
$x_{T}^{K}$ 41.0913$\pm$3.4621 38.3504$\pm$2.9553 0.9635$\pm$0.0656 0.9458$\pm$0.0678
$x_{z}^{K}$ 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
20% sampling steps $x_{T}$ 19.7481$\pm$7.8121 17.5994$\pm$7.7102 0.4664$\pm$0.2049 0.4312$\pm$0.1723
$x_{T}^{0}$ 38.4951$\pm$2.9590 36.4403$\pm$2.9405 0.9500$\pm$0.0600 0.9317$\pm$0.0656
$x_{T}^{K}$ 40.8341$\pm$3.8591 38.2336$\pm$3.0403 0.9631$\pm$0.0654 0.9421$\pm$0.0691
$x_{z}^{K}$ 40.9907$\pm$3.7419 38.3414$\pm$2.8912 0.9631$\pm$0.0657 0.9446$\pm$0.0682
pretrained with another dataset $x_{T}$ 25.5783$\pm$4.7741 24.2301$\pm$7.8034 0.4904$\pm$0.1506 0.4692$\pm$0.1549
$x_{T}^{0}$ 39.0112$\pm$2.7413 36.8603$\pm$2.7916 0.9510$\pm$0.0578 0.9351$\pm$0.6225
$x_{T}^{K}$ 40.9581$\pm$3.3235 38.2714$\pm$2.8405 0.9634$\pm$0.0663 0.9401$\pm$0.0702
$x_{z}^{K}$ 41.0920$\pm$3.2329 38.4204$\pm$2.9084 0.9636$\pm$0.6828 0.9432$\pm$0.0688
V-D Evaluation of the guided reconstruction process
To evaluate the improvements brought by the guided process, we plotted the following images: the image of the ground truth, zero-filled measurements, PGI ($x_{T}$), DGI ($x_{T}^{0}$), and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branches ($x_{T}^{K}$ and $x_{z}^{K}$) of an MR image from the test dataset. To mitigate potential influences from the sampling process, we designed the following three sampling methods:
  1. 1.
    original: the SMLD results are sampled using the pretrained score network, as described previously.
  2. 2.
    20% sampling steps: the SMLD results are sampled using the pretrained score network, but using only 20% of the original sampling steps.
  3. 3.
    pretrained with another dataset: we continued to train the score network on another brain dataset proposed in [69], and the images are sampled with the same hyperparameters in the original method.
The reconstructed images from the original sampling method at four-fold acceleration are shown in Fig. 5. The first row shows the corresponding images, and the second row shows the magnified details. The zero-filled image (second column) is blurred, and the PGI (third column) loses texture details. However, the artifacts in the DGI (fourth column) were effectively eliminated from the zero-filled image and PGI by DMs. Then, the GICs recovered the missing details in the DGIs and $x_{T}^{K}$ (fifth column) and $x_{z}^{K}$ (sixth column) presented satisfactory results when compared with the ground truth images (first column). The error maps in the third row exhibit that the massive errors in the zero-filled images and PGIs are obviously alleviated, the the image quality are further improved in $x_{T}^{K}$ and $x_{z}^{K}$. The reconstructed images with 20% sampling steps at four-fold acceleration are depicted in Fig. 6, which are also consistent with the observations in Fig. 5.
The quantitative results for the three sampling methods are presented in Table III. $x_{z}^{K}$, which is chosen as the final result, slightly outperforms $x_{T}^{K}$, and the other results obtained at four- or six-fold acceleration are consistent with the results observed from the images. Significantly reducing the sampling steps leads to further deterioration in the quality of SMLD results. However, the impact on the final reconstruction results is minimal, demonstrating the robustness of our proposed method to noise input, and the DMs are effective to extract valuable information from the noise input. We continue training the score network on another brain dataset to make perturbations on the pretrained model, which produces reconstructions of equally high quality through our workflow, demonstrating that our method imposes minimal requirements on the sampling process. In summary, this experiment vividly illustrates how our proposed workflow leverages SMLD results to extract crucial information, aiding in the reconstruction process and finally achieving high-quality results.
V-E Evaluation of the network robustness
Figure 7: PSNR (the first column) and SSIM (the second column) results of SGM-Net, MeDL-Net, and E2EVarNet with different amounts of training data at four- (the first row) and six-fold (the second row).
In this subsection we illustrate the network robustness with limited training data. Here, we compared E2E-VarNet and MeDL-Net, which have convincing performances and do not leverage guidance images, as shown in Table I. We also compared with three ablation networks (i.e. SG-MN-DM-US, SG-MN-MN-DG and SG-MN-US-DG). We reduced the number of training images of the different methods to different extents and compared their performances with our model. The results are presented in Fig. 7. We compared the PSNR and SSIM results of the above networks at four- and six-fold, respectively. We observed that the performances of the other networks deteriorated with decreasing training data. Meanwhile, the performance of SGM-Net was more stable even with only 20 training images. Compared with E2E-VarNet and MeDL-Net, our ablation networks also exhibit relatively better robustness. However, the degradation of image quality become more pronounced with the decrease of training data, which further demonstrating the necessity of the proposed modules. To sum up, this experiment demonstrates that proposed guidance process can effectively compensate for information in the missing training data and make the model robust.
VI Discussions
To make the sampling process robust and achieve better performance, most of the cutting-edge methods proposed to optimize the training of the generative models or the sampling steps. For example, the authors in [14] fed high-dimensional tensors to a noise conditional score network during training, and they designed homotopic iterations at the inference phase. The authors in [13] proposed a framework named score-POCS. During inference, the reconstruction process iterates between SDE solvers and data consistency steps. Compared with them, the most significant distinction in our workflow is that we bypass all the tedious hyperparameter tuning and network retraining steps and implement a naive sampling. The reconstruction performances are then not robust especially with distribution-shift test datasets, and the results are observed to be severely corrupted. We novelly treat the SMLD results as a generalized prior, and we design an unrolled network which can extract valuable information from the SMLD samples to guide reconstruction, while the components during sampling can also help to eliminate the artifacts in the network. Experimental results have demonstrated that compared with the naive sampling results, only small amounts of data are required to train the unrolled network and the performances can be greatly improved.
Another bottleneck in diffusion models is that the sampling process is relatively slow. To solve this problem, the authors in [70] proposed DiffuseRecon, and the reverse process can be guided with observed k-space signal. The authors designed a coarse-to-fine algorithm to save sampling time. In [71], the authors proposed deep ensemble denoisers and incorporated them with score-based generative models to reconstructed images. Consequently, the necessary iterations are reduced. In our workflow, we preserved the naive sampling process. However, the unrolled network have good tolerance to artifact-affected inputs, hence the sampling steps can also be great reduced.
In the proposed reconstruction workflow, the low-quality SMLD samples have been demonstrated to be effective to guide the network training and improve the final results. Meanwhile, the proposed workflow is also compatible to optimized SMLD steps with higher quality references, which show great potential to be exploited in future works. Besides, a limitation of the proposed method is that although the workflow exhibits considerable tolerance for the SMLD process, the unrolled network still requires some data for training to achieve superior performance. Training strategies such as self-supervising should be considered to further optimize the workflow in future works.
VII Conclusions
In this study, we proposed a novel workflow to eliminate the need for optimizing the forward and reverse process in the SMLD-based method for MRI reconstruction. Instead, an unrolled model-driven network is proposed to extract valuable information from naive sampling results. Subsequently, the SMLD results can be periodically updated and utilized to guide the network training process. The experimental results demonstrated that the proposed hybrid method outperforms the workflows solely based on diffusion models or unrolled networks. Additionally, our method alleviates the substantial requirements for sampling steps and training images. When compared with cutting-edge methods, our network consistently produces reconstructions of higher quality, showcasing its superiority in MRI reconstruction tasks.

### v2

@@
 Score matching with Langevin dynamics (SMLD) method has been successfully applied to accelerated MRI.
-However, the hyperparameters in the sampling process require subtle tuning, otherwise the results can be severely corrupted by hallucination artifacts, particularly with out-of-distribution test data.
-In this study, we propose a novel workflow in which SMLD results are regarded as additional priors to guide model-driven network training.
-First, we adopted a pretrained score network to obtain samples as preliminary guidance images (PGI) without the need for network retraining, parameter tuning and in-distribution test data.
-Although PGIs are corrupted by hallucination artifacts, we believe that they can provide extra information through effective denoising steps to facilitate reconstruction.
-Therefore, we designed a denoising module (DM) in the second step to improve the quality of PGIs.
-The features are extracted from the components of Langevin dynamics and the same score network with fine-tuning; hence, we can directly learn the artifact patterns.
-Third, we designed a model-driven network whose training is guided by denoised PGIs (DGIs).
-DGIs are densely connected with intermediate reconstructions in each cascade to enrich the features and are periodically updated to provide more accurate guidance.
-Our experiments on different sequences revealed that despite the low average quality of PGIs, the proposed workflow can effectively extract valuable information to guide the network training, even with severely reduced training data and sampling steps.
+However, the hyperparameters in the sampling process require subtle tuning, otherwise the results can be severely corrupted by hallucination artifacts, especially with out-of-distribution test data.
+To address the limitations, we proposed a novel workflow where naive SMLD samples serve as additional priors to guide model-driven network training.
+First, we adopted a pretrained score network to generate samples as preliminary guidance images (PGI), obviating the need for network retraining, parameter tuning and in-distribution test data.
+Although PGIs are corrupted by hallucination artifacts, we believe they can provide extra information through effective denoising steps to facilitate reconstruction.
+Therefore, we designed a denoising module (DM) in the second step to coarsely eliminate artifacts and noises from PGIs.
+The features are extracted from a score-based information extractor (SIE) and a cross-domain information extractor (CIE), which directly map to the noise patterns.
+Third, we designed a model-driven network guided by denoised PGIs (DGIs) to further recover fine details.
+DGIs are densely connected with intermediate reconstructions in each cascade to enrich the information and are periodically updated to provide more accurate guidance.
+Our experiments on different datasets reveal that despite the low average quality of PGIs, the proposed workflow can effectively extract valuable information to guide the network training, even with severely reduced training data and sampling steps.
 Our method outperforms other cutting-edge techniques by effectively mitigating hallucination artifacts, yielding robust and high-quality reconstruction results.
-Index Terms:
-MRI reconstruction, model-driven, diffusion model, compressed sensing, denoising score matching.
-I Introduction
-Magnetic Resonance Imaging (MRI) inherently requires long scanning times, leading to patient discomfort and degradation of image quality.
+keywords:
+MRI reconstruction , model-driven , diffusion model , compressed sensing , denoising score matching
+journal: Biomedical Signal Processing and Control
+[…]
+[…]
+1 Introduction
+Magnetic Resonance Imaging (MRI) inherently requires long scanning time, leading to patient discomfort and degradation of image quality.
 Undersampling the k-space data can proportionally reduce the scanning time but results in aliasing artifacts in the reconstructed images.
-The conventional methods for reconstructing high-quality images from a small subset of k-space data are mainly parallel imaging (PI) [1, 2] and compressed sensing (CS) [3, 4, 5].
+The conventional methods for reconstructing high-quality images from a small subset of k-space data are mainly based on parallel imaging (PI) [1, 2] and compressed sensing (CS) [3, 4, 5].
 In recent years, deep-learning (DL)-based methods have been successfully introduced in the field of MRI reconstruction and have shown convincing improvements over conventional methods [6, 7].
-For DL-based reconstruction, the proposed reconstruction models can be categorized into data-driven and model-driven unrolled networks.
-Part of the data-driven model was designed in an end-to-end manner to fit the mapping between the input measurements and reconstruction targets [8, 9, 10].
+Early networks are designed in an end-to-end manner to fit the mapping between the input measurements and reconstruction targets [8, 9].
 However, these methods still present problems such as lack of interpretability, and a relatively large dataset is required to fit the mapping.
-Recently, some data-driven networks have focused on distribution-learning-based reconstruction processes using diffusion models [11, 12, 13, 14, 15, 16, 17].
-Among them, denoising diffusion probabilistic models (DDPMs) [18, 19] and score matching [20, 21] with Langevin dynamics (SMLD) methods, which correspond to Varience Preserving (VP) Stochastic differential equations (SDEs) and Variaence Exploding (VE) SDEs, respectively [22], have shown great potential in solving inverse problems.
+Recently, some studies have focused on distribution-learning-based reconstruction processes using diffusion models [10, 11, 12, 13, 14, 15, 16].
+Among them, denoising diffusion probabilistic models (DDPMs) [17, 18] and score matching [19, 20] with Langevin dynamics (SMLD) methods, which correspond to variance preserving (VP) stochastic differential equations (SDEs) and variance exploding (VE) SDEs, respectively [21], have shown great potential in solving inverse problems.
 They have also been successfully extended to the MRI reconstruction field [22, 23, 21, 24].
@@
 The score network was trained on a brain MRI dataset.
-Images in odd columns are objective reality and those in even columns are sampling results.
-Model-driven unrolled networks are built in a different manner: conventional optimization algorithms are unrolled and realized in deep networks, and stacked cascades are used to mimic the iterative steps in the optimization procedure.
-Compared with data-driven methods, unrolled networks can be better interpreted as those in which conventional priors can be appropriately incorporated, and they have shown convincing results in recent years.
-Although stacking more cascades can improve the reconstruction performance, large-scale backbone networks can hardly be applied to unrolled networks, considering the limitations of GPU memory, and a trade-off must be established.
-However, data- and model-driven methods attempt to recover the same target images along different paths, and the complementarity between these methods should be considered.
-Moreover, we observed the convincing potential of SMLD methods for MRI reconstruction.
-Current score-matching networks require retraining when applied to different datasets and we observe artifacts on the sampled images without network retraining and subtle tuning of hyperparameters, as shown in Fig.
+Images in odd columns are ground truth and those in even columns are sampling results.
+Model-driven networks are built in a different manner: conventional optimization algorithms are unrolled and realized in deep networks, and stacked cascades are used to mimic the iterative steps in the optimization procedure.
+Compared with the other methods, unrolled networks offer better interpretability.
+Enhanced reconstruction performance is observed with learning based parameters and regularizers in unrolled networks compared with traditional optimization procedure.
+Although stacking more cascades can improve the reconstruction performance, large-scale backbone networks can hardly be applied to unrolled networks due to the limitations of GPU memory, and a trade-off must be established.
+Considering that distribution-learning-based and model-driven methods attempt to recover the same target images along different paths, the complementarity between these methods should be considered.
+Current score-matching networks require retraining to achieve superior results across varying datasets.
+We observed artifacts on the sampled images without network retraining and tuning of hyperparameters, as shown in Fig.
 1.
-However, we believe that the artifact-affected images can still provide information additional to the known undersampled measurements if the artifacts can be appropriately eliminated, thereby compensating for missing measurements.
-Instead of focusing on optimizing the forward and reverse process in SMLD methods, we aim to design a model-driven network which can extract valuable information from naive sampling process and achieve high-quality results robustly.
-We proposed a novel DL-based hybrid MRI reconstruction workflow, consisting of the following three main steps: (1) Sampling: Using known undersampled raw k-space data, we first obtain samples using SMLD methods.
+Despite these artifacts, we can observe that the artifact-affected images still provide information additional to the known undersampled measurements.
+Therefore, if the artifacts can be appropriately eliminated, naive samples from SMLD methods can compensate for missing measurements and provide complementary information to guide model-driven network training.
+To this end, instead of focusing on optimizing the forward and reverse process in SMLD methods, we aimed to design a model-driven network which can extract valuable information from naive sampling process and achieve high-quality results robustly.
+We proposed a novel DL-based hybrid MRI reconstruction workflow, consisting of the following three main steps: (1) Sampling: Using known undersampled raw k-space data, we first obtained initial samples using SMLD methods.
 We bypassed all the tuning steps in SMLD methods, such as tuning the hyperparameters or retraining the score networks on our dataset.
@@
 A PSN can be trained using out-of-distribution data.
-(2) Denoising: We believe that these preliminary reconstructions can provide additional information to guide the reconstruction process through appropriate denoising, leading to a better performance.
-Considering that the PGIs have relatively low quality and that we observe similar aliasing artifacts in the images, we designed a denoising module (DM) to remove the artifacts from the PGIs.
-We extracted features from the same score network after fine-tuning and concatenated them with the inputs in Langevin dynamics, which directly produces artifacts, to learn the noise pattern.
+(2) Denoising: We believe that these preliminary reconstructions can provide additional information to guide the reconstruction process through appropriate denoising, leading to better performance.
+Considering that the PGIs have relatively low quality and we observed similar artifacts in the images, we designed a denoising module (DM) to remove the artifacts and noises from the PGIs.
+We designed a score-based information extractor (SIE) and a cross-domain information extractor (CIE) to build mappings from extracted features to the noise patterns.
 Then, we used denoised PGIs (DGIs) as higher-quality guidelines for further reconstruction.
@@
 The DM and unrolled cascades were trained end-to-end.
-The experimental results show that although the quality of PGIs is even lower than that of the undersampled measurements, the artifacts were effectively eliminated by the DMs.
+Finally, the proposed workflow essentially reconstruct MR images in a coarse-to-fine manner.
+The denoising steps primarily restore low frequency patterns, while the guided reconstruction steps further recover fine details, achieving superior image quality.
+The experimental results show that although the average quality of PGIs is even lower than that of the undersampled measurements, the artifacts were effectively eliminated by the DMs.
 Meanwhile, the score network can not only help in the sampling process but also cooperate with the unrolled network and improve the denoising performance of the DMs.
-The proposed reconstruction workflow outperformed state-of-the-art methods through the guidance of DGIs.
+Finally, the proposed reconstruction workflow outperformed state-of-the-art methods through the guidance of DGIs.
 The main contributions of this study are summarized as follows:
-  1. 1.
-    We propose a novel DL-based MRI reconstruction workflow, and low-quality samples obtained using SMLD method are used as guidance during reconstruction.
-  2. 2.
-    We design an unrolled network, and the initially obtained samples are denoised and periodically updated to guide network training.
-  3. 3.
-    Experimental results show that without tuning the sampling steps or retraining the score network on target dataset, the low-quality samples can guide the proposed network to obtain robust and high-quality reconstructions, which outperforms the cutting-edge methods.
-II Related Works
-II-A Conventional approaches for accelerated MRI
-For the PI, multiple receiver coils are equipped to obtain raw k-space data simultaneously, and the missing data are recovered by exploiting redundancies among the data from different coils.
-Sensitivity encoding (SENSE)-type [26] methods work in the image domain, whereas others directly interpolate missing data points in k-space, such as simultaneous acquisition of spatial harmonics (SMASH) [27] and generalized autocalibration partial parallel acquisition (GRAPPA) [28].
+  • 1.
+    We propose a novel DL-based MRI reconstruction workflow, and low-quality samples obtained using naive SMLD method can be leveraged as guidance during reconstruction.
+  • 2.
+    We propose a DM with SIE and CIE blocks which can coarsely remove artifacts and noises in the SMLD samples, and an unrolled network is proposed to further recover fine details.
+  • 3.
+    Experimental results demonstrate that without tuning the sampling steps or retraining the score network on target dataset, the low-quality samples can guide the proposed network to obtain robust and high-quality reconstructions, which outperforms the cutting-edge methods.
+2 Related Works
+For PI methods, multiple receiver coils are equipped to obtain raw k-space data simultaneously, and the missing data are recovered by exploiting redundancies among the data from different coils.
+Sensitivity encoding (SENSE)-type [25] methods work in the image domain, whereas some others directly interpolate missing data points in k-space, such as simultaneous acquisition of spatial harmonics (SMASH) [26] and generalized autocalibration partial parallel acquisition (GRAPPA) [27].
 However, the reconstructions in PI approaches deteriorate at high acceleration factors.
@@
 However, conventional methods depend heavily on handcrafted procedures, and the iterative steps are time-consuming.
-Figure 2: (a) The workflow of the proposed SGM-Net. (b) The network structure of the sampling and denoising steps. (c) The network structure of the $k$th cascade.
-II-B Data-driven networks for MRI reconstruction
-For data-driven networks trained end-to-end, the mapping between the undersampled measurements and reconstructed images was established.
+For deep networks trained end-to-end, the mapping between the undersampled measurements and reconstructed images are established.
 For example, early studies adopted deep convolutional neural networks (CNNs) to reconstruct MR images from zero-filled inputs [34, 35].
@@
 In [37], a complex-valued deep residual network with a higher performance and smaller size was proposed.
-In addition, GAN-[39, 40]-based networks have shown convincing results [41, 42, 43].
-In addition to realizing image domain end-to-end learning, some studies have proposed direct learning mapping in k-space [44] or exploiting the complementarity between the image and k-space domains [9].
-Distribution learning is another category of data-driven networks.
+In addition to realizing image domain end-to-end learning, some studies have proposed direct learning mapping in k-space [38] or exploiting the complementarity between the image and k-space domains [9, 39].
+In [40], the authors proposed an unsupervised framework which learn from image and frequency domain in a multi-stage manner.
+Some other methods direct learned an image prior in generative models such as generative adversarial networks (GAN) [41, 42, 43, 44, 45, 46].
 For example, in [47] the authors used pre-trained generative models to present data distributions.
+More recently, diffusion models have emerged as new state-of-the-art generative models.
 In [22], the authors used score matching to estimate the data distribution, and samples were then generated using annealed Langevin dynamics.
 Subsequently, the authors proposed improved skills in [23] for higher stability and better performance, and exact log-likelihood computation was supported in [21].
-[25] first successfully applied CS-based generative models to overcome the MRI reconstruction problem and showed convincing results in out-of-distribution samples.
+[24] first successfully applied CS-based generative models to overcome the MRI reconstruction problem.
 In [15], the authors further optimized the sampling stage by automatically tuning the hyperparameters using Stein’s unbiased risk estimator (SURE).
@@
 Some works extended the idea of cold diffusion [48] to MRI reconstruction, and the process of adding Gaussian noise is replaced with undersampling operation [49].
-However, network retraining is necessary when changing the acquisition parameters, and large datasets are required to fit the mapping [48].
-II-C Unrolled networks for MRI reconstruction
-Unrolled networks, also known as model-driven networks, are constructed by unrolling iterative optimization steps into deep networks [49, 50, 51, 52, 53].
+However, these networks generally required large training datasets [50] and suffer from the lack of interpretability.
+Moreover, tuning the training and sampling process of diffusion models is non-trivial and crucial for achieving robust reconstruction outcomes.
+Unrolled networks, also known as model-driven networks, are constructed by unrolling iterative optimization steps into deep networks [51, 52, 53, 54, 55, 56, 57].
 For example, algorithms such as the alternating direction method of multipliers (ADMM) [51], iterative shrinkage-thresholding algorithm (ISTA) [52], variable splitting (VS) [54], and approximate message passing (AMP) have been successfully applied to solve the objective functions in MRI reconstruction.
@@
 The authors in [66] designed a deep low-rank and sparse network which explored 1D convolution during training.
-However, the complementarity between unrolled networks and data-driven methods has not been investigated; in this study, we propose a novel concept in which we treat naive sampling results from SMLD methods as guidance and unrolled cascades as denoising and refining processes.
-III Background
+In [67], ADMM was utilized to unroll a half-quadratic variable splitting based optimization process, and a dilated-convolution model was designed to predict the Lagrange multipliers.
+However, the complementarity between unrolled networks and distribution-learning-based methods has not been thoroughly investigated; in this study, we propose a novel concept in which we treat naive sampling results from SMLD methods as guidance and unrolled cascades as denoising and refining processes.
+3 Background
 In PI, the acquisition of multicoil measurements can be formulated as
@@
 The expanding operator $\mathcal{E}(\cdot)$ calculates the coil-specific images from $x$, defined as
-$\mathcal{E}(x)=(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...),i=1,2,...,C$ (3)
+$\mathcal{E}(x)=(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...\mathcal{S}_{C}x)=(x_{1},x_{2},...,x_{C}),i=1,2,...,C$ (3)
 where $\mathcal{S}_{i}$ is the SM of the $i$th coil. Conversely, the reduced operator $\mathcal{R}$ integrates the coil-specific images into a single image defined as follows:
-$\mathcal{R}(\mathcal{S}_{1}x,\mathcal{S}_{2}x,...)=\sum_{i=1}^{C}\mathcal{S}_{i}x$ (4)
+$\mathcal{R}(x_{1},x_{2},...,x_{C})=\sum_{i=1}^{C}\mathcal{S}_{i}^{*}x_{i}$ (4)
 and the backward operator is expressed as
 $\mathcal{A}^{*}(\cdot)=\mathcal{R}\circ\mathcal{F}^{-1}\circ\mathcal{P}(\cdot)$ (5)
-It is ill-posed to recover missing data directly from $y$. Hence, a regularization term $\lambda\mathcal{U}(x)$ is routinely added to narrow the solution space in the CS, which is
-$\hat{x}=\mathop{\arg\min}\limits_{x}\sum_{i=1}^{C}\|\mathcal{P}\mathcal{F}\mathcal{S}_{i}x-y\|_{2}^{2}+\lambda\mathcal{U}(x)$ (6)
-For example, $\lambda\Psi(x)$ is a common regularization term that imposes the sparsity constraint. Then, gradient descent can be used to solve the problem iteratively as follows:
+It is ill-posed to recover missing data directly from $y$. Hence, a regularization term $\lambda\mathcal{U}(x)$ is routinely added to narrow the solution space, which is
+$\hat{x}=\mathop{\arg\min}\limits_{x}\|\mathcal{P}\mathcal{F}\mathcal{E}x-y\|_{2}^{2}+\lambda\mathcal{U}(x)$ (6)
+For example, $\lambda\Psi(x)$ is a common regularization term that imposes the sparsity constraint in CS-MRI. Then, gradient descent can be used to solve the problem iteratively as follows:
 $x^{t+1}=x^{t}-\alpha^{t}(\mathcal{A}^{*}(\mathcal{A}x^{t}-y)+\lambda\psi(x^{t}))$ (7)
-where $\alpha^{t}$ is the step size in the $t$th iteration and $\psi(x^{t})$ is the gradient of $\Psi(x)$.
-IV Methods
+where $\alpha^{t}$ is the step size in the $t$th iteration and $\psi(x^{t})$ is the gradient of $\Psi(x^{t})$.
+4 Methods
+Figure 2: (a) The workflow of the proposed SGM-Net. (b) The network structure of the denoising module and the SIE and CIE blocks. (c) The network structure of the $(k+1)$th GIC.
 We proposed a score-based generative priors guided model-driven Network (SGM-Net), which is illustrated in Fig.
@@
 2 (a)) were first obtained using Langevin dynamics and then used as PGIs for the follow-up steps.
-The gradients of the data distribution were obtained using PSN.
+The gradients of the data distribution were obtained using a PSN.
 Instead of carefully training and tuning the SMLD process to obtain high-quality samples, we sampled $x_{T}$ from an out-of-distribution pretrained network with fitted hyperparameters.
-Then, we treat the artifacts in $x_{T}$ as removable and propose a DM to eliminate noise and enhance $x_{T}$.
-The information available from the sampling steps was integrated and fed into the DM, as shown in Fig.
+Then, we treated the artifacts in $x_{T}$ as removable and proposed a DM to enhance $x_{T}$.
+We designed a SIE and a CIE to extract features in DMs, as shown in Fig.
 2 (b).
-Finally, DGIs were used to guide an unrolled network, as shown in Fig.
+Finally, DGIs were used to guide an unrolled network training, as shown in Fig.
 2 (c).
 Details of the proposed method are presented in the following subsections.
-IV-A Score matching and Langevin dynamics
-The first step of the proposed workflow is to obtain PGIs with SMLD to guide the subsequent steps.
-In this study, we adopted the PSN proposed in [23], which was successfully extended to MRI reconstruction in [25].
-To improve the generality of our workflow, the score network was pretrained with out-of-distribution data and tedious tuning and retraining steps were bypassed.
-The network, remarked as $s_{\theta}$, is a four-cascade RefineNet [63] with dilated convolutions and instance normalization, which is used to approach the score function of $p(x)$ (that is, $\bigtriangledown_{x}\log p(x)$) and we have $s_{\theta}\approx\bigtriangledown_{x}\log p(x)$.
-Average pooling layers were adopted to replace all maximum pooling layers in RefineNet, and ELU was adopted as the activation function.
-Then, the annealed Langevin method-based sampling process in MRI can be represented as
+4.1 Score matching and Langevin dynamics
+The first step of the proposed workflow is to obtain PGIs using SMLD method.
+In this study, we adopted a PSN proposed in [22], which was successfully extended to MRI reconstruction in [24].
+To improve the generality of our workflow, the score network was pretrained with out-of-distribution data, and the tedious tuning and retraining steps were bypassed.
+The network, remarked as $s_{\theta}$, is a RefineNet [68] with dilated convolutions and conditional instance normalization, which is used to approach the score function of $p(x)$ (that is, $\bigtriangledown_{x}\log p(x)$) and $s_{\theta}\approx\bigtriangledown_{x}\log p(x)$.
+$s_{\theta}$ is a six-layer U-shaped network.
+The encoders are designed based on ResNet blocks and the decoders are based on RefineNet blocks.
+The annealed Langevin method-based sampling process in MRI can be represented as
 $x_{t+1}=x_{t}+\eta_{t}(s_{\theta}(x_{t};\sigma_{t})+\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{t})}{\gamma_{t}^{2}+\sigma^{2}})+\sqrt{2\eta_{t}}\xi_{t}$ (8)
-where $\xi_{t}\sim\mathcal{N}(0,\textit{I})$ and $t=0,1,2,..,T$.
+where $\xi_{t}\sim\mathcal{N}(0,\textit{I})$ and $t=0,1,2,..,T-1$.
 $\eta_{t}$ is the step size in the $t$th iteration and $\eta_{t}=\epsilon\cdot\frac{\sigma^{2}_{j}}{\sigma^{2}_{L}}$.
@@
 $\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{t})}{\gamma_{t}^{2}+\sigma^{2}}$ is the consistency term introduced to exploit known measurements.
-Finally, $x_{T}$ approaches the target as $T$ becomes sufficiently large, which is used as a guide in the follow-up steps.
-IV-B Denoising module
-Over-smoothing details and hallucination artifacts were observed on the PGIs obtained in the first step, and the denoising step aimed to initially eliminate artifacts and compensate for missing details.
-First, we designed a KI-branch to learn an denoising mapping from the data between the frequency domain and the image domain.
-Considering that artifacts usually have different frequencies, which makes it possible to separate them in the k-space, we transferred $x_{T}$ back to the k-space and used a U-Net with a skip connection to learn the mapping to the denoised outputs.
-Subsequently, U-Net learned to eliminate the noisy signals in k-space.
-A data consistency (DC) block was concatenated to the U-Net to maintain consistency with the measurements.
-In addition, the complementarity between the k-space and image domain could also help obtain better performance.
-We then used an inverse Fourier transform to transfer the k-space data back to the image domain, which is denoted as $x_{d}$, as shown in Fig.
-2 (b).
-Then, we designed an input-enriched U-Net, marked as $N_{d}$, to further eliminate the noises in $x_{d}$.
-Fundamentally, the score network is used to generate clean images from noises and the artifacts are generated during the sampling process.
-Consequently, there could exist an underlying mapping between the artifacts and the elements involved in the sampling process (i.e.
-equation (8)).
-Therefore, with the input $x_{T}$ in the DM, we fed $x_{T}$, $s_{\theta}(x_{T};\sigma_{T}$) and $\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{T})}{\gamma_{T}^{2}+\sigma^{2}}$ as the input elements.
-We fine-tuned $s_{\theta}$, marked as $f_{s_{\theta}}$ that we removed the last batch normalization and activation layers to avoid undesirable loss of information, and concatenated a fine-tuning convolutional (FTC) layer to learn the noise pattern from the features.
-Only the FTC layer in $f_{s_{\theta}}$ was updated during training.
-The parameters in the other parts of $f_{s_{\theta}}$ were fitted, and the Langevin sampling results were fed to $f_{s_{\theta}}$ such that the FTC layers could directly learn the noise pattern from the score-matching process for each image.
-Besides, to avoid newly introduced noises when obtaining $x_{d}$ from $x_{T}$, $x_{d}$ was also fed to $N_{d}$.
-Therefore, the enriched input features of $N_{d}$ have four components.
-By concatenating and feeding the four components into the DM, the denoising process is denoted as:
-$x_{T}^{0}=x_{d}+N_{d}\{x_{p},x_{T},\frac{\mathcal{A}^{*}(y-\mathcal{A}x_{T})}{\gamma_{T}^{2}+\sigma_{T}^{2}},x_{d}\}$ (9)
+Finally, $x_{T}$ approaches the target as $T$ becomes sufficiently large, which is used as a guidance in the follow-up steps.
+4.2 Denoising module
+The PGIs generated in the first step exhibits over-smooth details and hallucination artifacts, and the denoising step aims to coarsely eliminate the noises and artifacts in PGIs.
+To achieve this, we designed a CIE and a SIE block to extract features, which were then used within a residual structure to generate denoised outputs.
+Consequently, the mappings between extracted features and the noise patterns were established.
+The CIE blocks were designed to exploit the complementarity between the frequency domain and the image domain.
+As depicted in Fig 2 (b), first we used the expand operator to obtain the corresponding k-space data of $x_{T}$, marked as $y_{T}$.
+Then, $y_{T}$ was reshaped and the coil channel was combined with the feature channel before fed to a U-Net.
+In this study, all complex-valued MR images were stored in two distinct channels during the network training.
+Therefore, the number of input and output channels of the U-Net was set to 30 (with 15-coil data).
+A data consistency (DC) block was concatenated to the output of U-Net, marked as $y_{c}$, to maintain consistency with the measurements, and the $j$th coefficient in $y_{c}$ was updated as:
+$y_{d}(j)=D(y_{c}(j))=\left\{\begin{array}[]{l}y_{c}(j),\ \ \ \ \ \ \ \ \ j\notin\Omega\\ \frac{y_{c}(j)+\mu y(j)}{1+\mu},\ \ \ \ j\in\Omega\\ \end{array}\right.$ (9)
+where $\Omega$ is the subset of sampled points from the complete k-space data.
+$y_{c}$ and $y_{d}$ are the input and output of the DC blocks, respectively, and $y$ is the undersampled measurements.
+We then used an reduce operator to transfer the k-space data back to the image domain, which is denoted as $x_{d}$.
+The CIE blocks could extract the features associated with the noise patterns directly in k-space.
+Considering the other components in DMs operate in the image domain, the output $x_{d}$ is expected to compensate for the missing information in the other blocks, thereby enhancing overall performance.
+Essentially, score network is used to generate clean images from noise, and the artifacts can be considered as generated during the sampling process.
+Consequently, there could exist an underlying mapping between the artifacts and the elements involved in the score network.
+Therefore, we designed a SIE to extract features correlated to the noise patterns from the PSN.
+As depicted in Fig 2 (b), the PSN is a six-layer RefineNet and the features in deeper RefineNet blocks (RBs) closely associated with the score of the inputs, which could potentially compromise noise-related information.
+Moreover, considering that the optimal level of features for learning the noise is unrevealed, we fine-tuned the PSN and extract features across multiple layers.
+Specifically, to build the mapping to the noise in $x_{T}$, we directly fed $x_{T}$ to PSN and during training, all the parameters in the PSN was fitted.
+Then, we concatenated convolutional blocks to the last four RBs in PSN to learn multi-level features.
+Each convolutions was followed by a batch normalization and a ReLU function.
+Then, deconvolutions were used to double the feature sizes, and the features were concatenated with the counterparts extracted in the shallower RBs.
+Each convolutions and deconvolutions was followed by a batch normalization block and a ReLU function.
+$1\times 1$ convolutions were first applied to the output of RBs to adjust the number of channels, thereby constraining the module scale.
+As a results, different levels of features can be simultaneously captured to maximize the stability and accuracy of the mapping established to noise.
+The output of the SIE is marked as $x_{p}$, which was concatenated with $x_{d}$ and fed to a U-Net, marked as $N_{d}$, to learn the noise in $x_{T}$.
+Besides, to avoid newly introduced noises when obtaining $x_{d}$ and $x_{p}$ from $x_{T}$, $x_{T}$ was also fed to $N_{d}$.
+Therefore, the enriched input features of $N_{d}$ have three components.
+By concatenating and feeding them into the $N_{d}$, the denoising process is denoted as:
+$x_{T}^{0}=x_{T}+N_{d}\{x_{T},x_{p},x_{d}\}$ (10)
 […]
-Figure 3: Examples of ground truth images (the first column), undersampled images (the second column) and the corresponding k-space trajectories (the third column).
-The images from the first and second rows are selected from the Knee dataset with four-fold and six-fold random Cartesian mask.
-The images from the third and fourth row are selected from the fastMRI dataset with four-fold and six-fold equispaced Cartesian mask.
+4.3 Unrolled network with PGI-guided densely-connected cascades
+The DM aims to realize a coarse denoising, and the third step of the proposed workflow is to further improve the image quality and recover the fine details from denoised guidance images (DGIs) (that is, $x_{T}^{0}$).
+The undersampled measurements $x_{z}^{0}$ and $x_{T}^{0}$ were used as two different known information during unrolled network training.
+As shown in Fig.
+2, the network consists of several DGI-guided densely connected cascades (GICs).
+The inputs of the $(k+1)$th GICs are $x_{T}^{k}$ and $x_{z}^{k}$ where $k=0,1,...,K-1$, and $x_{z}^{0}$ is the coil-combined zero-filled image of the undersampled measurements.
+In each GIC, $x_{z}^{k}$ were periodically updated to approach the target images, whereas the DGIs $x_{T}^{k}$ were simultaneously updated to provide more accurate guidance, forming a parallel network structure.
+$I$ Data fidelity (DF) and $I$ U-Net blocks were alternately stacked in each GIC to solve (7), and U-Net blocks served as learning based regularizers.
+For the $x_{T}^{k}$ branch, the calculations in DF blocks can be represented as
+$r_{T}^{k_{i+1}}=x_{T}^{k_{i}}-\alpha^{k_{i}}(\mathcal{A}^{*}(\mathcal{A}x_{T}^{{k_{i}}}-y)),i=0,1,...,I-1$ (11)
+where $x_{T}^{k_{0}}$ is equal to $x_{T}^{k}$. For the $x_{z}^{k}$-branch, the DF blocks are given by
+$r_{z}^{k_{i+1}}=x_{z}^{k_{i}}-\alpha^{k_{i}}(\mathcal{A}^{*}(\mathcal{A}x_{z}^{k_{i}}-y)),i=0,1,...,I-1$ (12)
+where $x_{z}^{k_{0}}$ is equal to $x_{z}^{k}$.
+In addition, we proposed a densely connected structure to enrich the input information for each U-Net block.
+This is based on the idea that DL-based methods are more flexible than the conventional algorithms, allowing for learning from richer inputs across various stages and extracting more valuable features.
+Consequently, the densely connected information can mitigate the impact of inaccurate cascades and establish robust mappings to the targets.
+For the $x_{T}^{k}$-branch, the initial input and each output of DF blocks corresponding to a U-Net block were concatenated before being fed into U-Net.
+Therefore, we have:
+$x_{T}^{k_{i+1}}=U_{T}^{i+1}\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}}\},i=0,1,...,I-1$ (13)
+where $U_{T}^{i+1}$ denotes the $(i+1)$th U-Net block in the $x_{T}^{k}$-branch, $x_{T}^{k}$ denotes the input to the corresponding branch, and $r_{T}^{k_{i}}$ denotes the output of the $i$th DF block.
+A similar strategy was implemented in the $x_{z}^{k}$-branch.
+Considering that we take the final output of the $x_{z}^{k}$ branch in the last GIC as the final reconstruction, the $x_{z}^{k}$ branch is essentially the main branch, and the $x_{T}^{k}$ branch is simultaneously used to guide the reconstruction in the main branch.
+Therefore, the concatenated information from the $x_{T}^{k}$ branch was also fed to the $x_{z}^{k}$ branch to learn from the features in different stages, represented as
+$x_{z}^{k_{i+1}}=U_{z}^{i+1}\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\},i=0,1,...,I-1$ (14)
+After the DF and U-Net blocks, we designed a skip connection and a DC block at the end of the $x_{T}^{k}$ branch, and the produced $x_{T}^{k+1}$ is fed to the next GIC as one of the two inputs as follows:
+$x_{T}^{k+1}=D(x_{T}^{k}+x_{T}^{k_{I}})$ (15)
+For $x_{z}^{k}$, we further incorporated a four-layer attention module (AM), marked as $\mathcal{M}_{A}$, between the last U-Net and skip connection, as shown in Fig 2 (c).
+Information fed to the last U-Nets in the two branches were concatenated and simultaneously fed to the AM to produce an attention map
+$m_{A}=\mathcal{M}_{A}\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{I}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{I}}\}$ (16)
+Through a skip connection and DC block, $x_{z}^{k+1}$ is given as
+$x_{z}^{k+1}=D(x_{z}^{k}+x_{z}^{k_{I}}\odot m_{A}+x_{T}^{k_{I}}\odot((\textbf{J}-m_{A}))$ (17)
+where J is an all-ones matrix and $\odot$ denotes the Hadamard product.
+The element-wise attention results in a deeper fusion between the two branches and optimizes the intermediate results in each $x_{z}^{k}$ branch, which amplifies the likelihood for subsequent GICs to converge towards the targets.
+4.4 loss functions
+In this study, the overall loss for network training was designed as:
+$\mathcal{L}{(x_{T}^{0},x_{g})}+\mathcal{L}{(x_{z}^{K},x_{g})}+\sum_{k=1}^{K}w_{k}[\mathcal{L}{(x_{T}^{k},x_{g})}+\mathcal{L}{(x_{z}^{k},x_{g})}]$ (18)
+where $x_{g}$ is the target image.
+$x_{T}^{0}$ amd $x_{z}^{K}$ are the final outputs of DM and the unrolled network.
+$x_{T}^{k}$ and $x_{z}^{k}$ are intermediate results from the two branches in the unrolled network.
+$\{w_{k}\}_{k=1}^{K}$ is a sequence assigns heavier weights to the intermediate results from deeper GICs and $w_{k}=10^{\frac{k-K}{K-1}}$ [67].
+When training on the coronal PD sequence and FastMRI dataset, we adopted mean square error as the loss function that $\mathcal{L}:=\mathcal{L}_{MSE}$.
+For the coronal PDFS sequence, we further included a SSIM loss that $\mathcal{L}:=\mathcal{L}_{MSE}+\mathcal{L}_{SSIM}$ which equally contribute with the MSE loss during training to improve reconstruction outcomes.
+4.5 Implementation details
+We adopted the score network $s_{\theta}$ released by the authors in [24], which was pretrained on FastMRI brain dataset.
+We followed the proposed hyperparameters in that manuscript [^1] https://github.com/utcsilab/csgm-mri-langevin during sampling, and we did not retrain the network on our knee dataset.
+In SGM-Net, $K$ was set to five and $I$ was set to three.
+The numbers of filters in the U-Nets from DMs were set to 32, 64, 128, and 256.
+The numbers of filters in all U-Nets from GICs were set to 16, 32, 64, and 128, respectively.
+In SIE blocks, the number of output channels of the RBs were set to 128, 128, 256, 256, 256 and 512.
+The number of channels of the inputs and outputs in $1\times 1$ convolutions, which are concatenated to RBs, were set to (128, 16), (128, 32), (256, 64) and (256, 128).
+For the last $1\times 1$ convolution, the number of input and output channels were 16 and 2, respectively.
+The CBRs and DBRs halved the number of channels of their inputs.
+The step size in DF blocks and the $\mu$s in DC blocks were learned directly in the network.
+The proposed SGM-Net was trained on a UBUNTU 20.04 LTS system with a 3090 GPU (24GB) for 100 epoches, and the learning rate was set to 1e-3.
+The batch size was set to 1.
+5 Experiments and Results
+5.1 Dataset
+Figure 3: Examples of ground truth images (the first column), undersampled images (the second column) and the corresponding k-space trajectories (the third column) from different datasets.
+In this study, we adopted a publicly available 15-coil NYU knee dataset [53] to train our model and compared it with other methods.
+The knee dataset consisted of five 2D turbo spin-echo (TSE) sequences, each of which contained volumes scanned using a 3T scanner from 20 patients.
+Each volume consists of approximately 40 slices.
+We conducted experiments on coronal proton density (PD; TR=2750 ms, TE=27 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 35-42 slices), and coronal fat-saturated PD (PDFS; TR=2870 ms, TE=33 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 33-44 slices) sequences.
+The coil sensitivity maps were precomputed using ESPRIRiT and included in the dataset.
+The training, validation, and testing sets were randomly divided at a ratio of 3:1:1.
+Random Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration.
+8% lines at the center region are preserved.
+FastMRI [8] is a publicly available dataset for accelerated MRI, which has widely been adopted in relevant studies.
+Considering the limitations of our hardware, we randomly selected a subset of the 15-coil knee dataset to evaluate our model and demonstrate its effectiveness.
+The data were acquired using the 2D TSE protocol, containing PD and PDFS sequence scanning from 3T or 1.5T systems (TR=2200-3000 ms, TE=27-34 ms, in-plane resolution 0.5$\times$0.5 mm2, slice thickness 3 mm).
+A total of 419, 33, and 375 slices were used for training, validation, and testing sets, respectively.
+Coil sensitivity maps are not available in the dataset, and related studies have been conducted to estimate and optimize sensitivity maps during training.
+However, we did not focus on SM estimation in this study and used pretrained SMs in our model to highlight the improvements brought about by the proposed methods.
+In the FastMRI dataset, we calculated the SMs using the sigpy [69] toolbox.
+Equispaced Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration.
+8% lines at the center region are preserved.
+Examples of training images and the corresponding k-space sampling trajectories are shown in Fig.
+3.
+5.2 comparisons with the cutting-edge methods
+Table 1: The quantitative results on Coronal PD and PDFS sequences, as well as the FastMRI dataset, compared with those of cutting-edge methods.
+Sequence Method PSNR SSIM
+Coronal PD zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
+TV 33.7498$\pm$3.1753 32.3290$\pm$3.2159 0.8851$\pm$0.0657 0.8658$\pm$0.0725
+U-Net 36.8984$\pm$2.7072 34.7352$\pm$2.7863 0.9361$\pm$0.0610 0.9133$\pm$0.0682
+D5C5 38.7903$\pm$2.8408 35.9697$\pm$2.7990 0.9509$\pm$0.0587 0.9259$\pm$0.0689
+ISTA-Net 39.2910$\pm$2.8423 35.5220$\pm$3.0915 0.9471$\pm$0.0567 0.9138$\pm$0.0717
+VS-Net 40.1649$\pm$2.8767 37.0604$\pm$2.7927 0.9591$\pm$0.0569 0.9358$\pm$0.0648
+E2E-VarNet 39.3592$\pm$2.9368 36.8949$\pm$3.3872 0.9542$\pm$0.0662 0.9342$\pm$0.0775
+ReVarNet 40.2545$\pm$2.8766 37.3683$\pm$2.7869 0.9590$\pm$0.0588 0.9379$\pm$0.0677
+MeDL-Net 40.9295$\pm$2.9072 38.1052$\pm$2.7622 0.9614$\pm$0.0557 0.9417$\pm$0.0663
+vSharp 40.3121$\pm$2.7773 37.5657$\pm$2.5346 0.9341$\pm$0.0575 0.8982$\pm$0.0660
+SGM-Net (ours) 41.1981$\pm$3.3802 38.4491$\pm$3.2166 0.9634$\pm$0.0682 0.9452$\pm$0.0790
+Coronal PDFS zero-filled 32.0299$\pm$2.0273 31.2221$\pm$2.0575 0.7964$\pm$0.0998 0.7587$\pm$0.1172
+TV 33.7256$\pm$1.9058 32.7328$\pm$1.8850 0.8124$\pm$0.0752 0.7920$\pm$0.0796
+U-Net 34.4367$\pm$0.6572 33.2028$\pm$2.5662 0.8183$\pm$0.0106 0.7758$\pm$0.1222
+D5C5 34.9671$\pm$2.8261 33.4189$\pm$2.4765 0.8250$\pm$0.1106 0.7854$\pm$0.1223
+ISTA-Net 34.4990$\pm$2.6389 33.4569$\pm$2.5758 0.8198$\pm$0.1066 0.7854$\pm$0.1214
+VS-Net 34.6847$\pm$2.7134 33.5010$\pm$2.6100 0.8259$\pm$0.1073 0.7842$\pm$0.1229
+E2E-VarNet 34.6274$\pm$2.7331 33.1626$\pm$2.4742 0.8234$\pm$0.1101 0.7804$\pm$0.1244
+ReVarNet 34.8522$\pm$2.7461 33.3146$\pm$2.5604 0.8272$\pm$0.1080 0.7832$\pm$0.1232
+MeDL-Net 35.5931$\pm$2.7307 34.1409$\pm$2.5501 0.8349$\pm$0.1025 0.7922$\pm$0.1255
+vSharp 35.5536$\pm$2.6244 34.3442$\pm$2.4485 0.8495$\pm$0.0940 0.8210$\pm$0.0990
+SGM-Net (ours) 36.1426$\pm$2.6821 34.7450$\pm$2.6737 0.8648$\pm$0.0946 0.8395$\pm$0.0987
+FastMRI zero-filled 33.2828$\pm$3.0594 32.5069$\pm$3.0657 0.8857$\pm$0.0709 0.8680$\pm$0.0791
+TV 35.5512$\pm$2.9189 33.8441$\pm$2.8703 0.8884$\pm$0.0635 0.0201$\pm$0.0152
+U-Net 35.2158$\pm$2.5925 33.7864$\pm$2.5331 0.8562$\pm$0.0574 0.8091$\pm$0.0637
+D5C5 36.4550$\pm$3.0246 34.5636$\pm$2.8510 0.9113$\pm$0.0649 0.8849$\pm$0.0749
+ISTA-Net 36.5753$\pm$2.9887 34.6229$\pm$2.8269 0.9134$\pm$0.0656 0.8837$\pm$0.0752
+VS-Net 37.3042$\pm$3.1130 35.5335$\pm$2.8513 0.9202$\pm$0.0637 0.8916$\pm$0.0736
+E2E-VarNet 37.5951$\pm$3.0129 35.6696$\pm$2.8326 0.9217$\pm$0.0635 0.8978$\pm$0.0726
+ReVarNet 37.7652$\pm$3.0632 35.7778$\pm$2.6971 0.9295$\pm$0.0589 0.9084$\pm$0.0675
+MeDL-Net 37.6791$\pm$2.9644 35.9594$\pm$2.7401 0.8984$\pm$0.0643 0.8882$\pm$0.0717
+vSHARP 38.0836$\pm$3.1283 35.7326$\pm$2.8239 0.9134$\pm$0.0645 0.8621$\pm$0.0917
+SGM-Net (ours) 38.8241$\pm$3.3808 36.8730$\pm$3.0364 0.9344$\pm$0.0588 0.9094$\pm$0.0707
 coronal PD 4$\times$ acceleration
@@
 Figure 4: Examples of the reconstructed images of different methods from coronal PD sequence and FastMRI dataset at 4$\times$ and 6$\times$ acceleration.
-The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps corresponding to the objective reality.
+The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps corresponding to the ground truth images.
 The methods from left to right: 1.
-objective reality; 2.
+ground truth; 2.
 zero-filled; 3.
@@
 MeDL-Net; 11.
+vSHARP; 12.
 SGM-Net (ours).
-IV-C Unrolled network with PGI-guided densely-connected cascades
-The third step of the proposed workflow is to design an unrolled network to obtain higher-quality reconstructions from denoised guidance images (DGIs) (that is, $x_{T}^{0}$), and the undersampled measurements $x_{z}^{0}$ and $x_{T}^{0}$ are used as two different constraints during unrolled network training.
-As shown in Fig.
-2, the network consists of several DGI-guided densely connected cascades (GICs).
-The inputs of GICs are $x_{T}^{k}$ and $x_{z}^{k}$ where $k=0,1,...,K$, and $x_{z}^{0}$ is the coil-combined zero-filled image of the undersampled measurements.
-In each GIC, the undersampled measurements $x_{z}^{k}$ were periodically updated to approach the target images, whereas the DGIs $x_{T}^{k}$ were simultaneously updated to provide more accurate guidance, forming a parallel network structure.
-$I$ Data fidelity (DF) and $I$ U-Net blocks were alternately stacked in each GIC to solve (7), and U-Net blocks were used to learn the gradient of $\Psi(x)$ from the data.
-For the $x_{T}^{k}$ branch, the calculations in DF blocks can be represented as
-$r_{T}^{k_{i+1}}=x_{T}^{k_{i}}-\alpha^{k_{i}}(\mathcal{A}^{*}(\mathcal{A}x_{T}^{{k_{i}}}-y)),i=0,1,...,I-1$ (10)
-where $x_{T}^{k_{0}}$ is equal to $x_{T}^{k}$. For the $x_{z}^{k}$-branch, the DF blocks are given by
-$r_{z}^{k_{i+1}}=x_{z}^{k_{i}}-\alpha^{k_{i}}(\mathcal{A}^{*}(\mathcal{A}x_{z}^{k_{i}}-y)),i=0,1,...,I-1$ (11)
-where $x_{z}^{k_{0}}$ is equal to $x_{z}^{k}$.
-In addition, we propose a densely connected structure to enrich the input information for each U-Net block.
-For the $x_{T}^{k}$-branch, the initial input and each output of DF blocks corresponding to a U-Net block are concatenated before being fed into U-Net.
-Therefore, we have:
-$x_{T}^{k_{i+1}}=U_{T}^{i}(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}}\})$ (12)
-where $U_{T}^{i}$ denotes the $i$th U-Net block in the $x_{T}^{k}$-branch, $x_{T}^{k}$ denotes the input to the corresponding branch, and $r_{T}^{k_{i}}$ denotes the output of the $i$th DF block.
-A similar strategy is implemented in the $x_{z}^{k}$-branch.
-Considering that we take the final output of the $x_{z}^{k}$ branch in the last GIC as the final reconstruction, the $x_{z}^{k}$ branch is essentially the main branch, and the $x_{T}^{k}$ branch is simultaneously used to guide the reconstruction in the main branch.
-Therefore, the concatenated information from the $x_{T}^{k}$ branch is also fed to the $x_{z}^{k}$ branch to learn from the features from different stages, represented as
-$x_{z}^{k_{i+1}}=U_{z}^{i}(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\})$ (13)
-After the DF and U-Net blocks, we designed a skip connection and a DC block at the end of the $x_{T}^{k}$ branch, and the produced $x_{T}^{k+1}$ is fed to the next GIC as one of the two inputs as follows:
-$x_{T}^{k+1}=DC(x_{T}^{k}+x_{T}^{k_{I}})$ (14)
-For $x_{z}^{k}$, we further adopted a four-layer attention module (AM) [64] between the last U-Net and skip connection.
-The information fed to the last DF block is concatenated and simultaneously fed to the AM to produce an attention map
-$M_{A}=AM(\{x_{T}^{k},r_{T}^{k_{1}},r_{T}^{k_{2}},...,r_{T}^{k_{i+1}},x_{z}^{k},r_{z}^{k_{1}},r_{z}^{k_{2}},...,r_{z}^{k_{i+1}}\})$ (15)
-Through a skip connection and DC block, $x_{z}^{k+1}$ is given as
-$x_{z}^{k+1}=DC(x_{z}^{k}+x_{z}^{k_{I}}\odot M_{A}+x_{T}^{k_{I}}\odot((\textbf{J}-M_{A}))$ (16)
-where J is an all-ones matrix and $\odot$ denotes the Hadamard product.
-Element-wise attention results in a deeper fusion between the two branches and optimizes the intermediate results in each $x_{z}^{k}$ branch.
-IV-D loss functions
-In this study, the loss function for unrolled network training was designed as
-$Loss=\gamma_{1}(x_{d}+x_{T}^{0}+x_{z}^{K})+\gamma_{2}(\sum_{k=1}^{K}x_{T}^{k}+\sum_{k=1}^{K}x_{z}^{k})$ (17)
-where $x_{d}$ amd $x_{T}^{0}$ are important guidelines for ensuring reconstruction performance, and $x_{z}^{K}$ is the result.
-$\sum_{k=1}^{K}x_{T}^{k}$ and $\sum_{k=1}^{K}x_{z}^{k}$ are intermediate results.
-We set $\gamma_{1}=1$ and $\gamma_{2}=0.1$ to distribute more weights for the guidance and results and fewer weights to help the network converge.
-IV-E Training details
-We adopted the score network $s_{\theta}$ released by the authors in [25], which is pretrained on a brain dataset.
-We followed the proposed hyperparameters in that manuscript [^1] https://github.com/utcsilab/csgm-mri-langevin during sampling, and we do not retrain the network on our knee dataset.
-The numbers of the four-cascade filters were 128, 256, 512, and 1024.
-The number of filters in the FTC layer is set to 2.
-In SGM-Net, $K$ is set to five and $I$ is set to three.
-The numbers of filters in all U-Nets in the DM and GICs were set to 16, 32, 64, and 128.
-The step size in DF blocks was learned directly in the network.
-All complex-valued MR images were stored in two channels during the network training.
-In particular, the first U-Net in DM uses k-space data as input, and the coil channel is combined with the feature channel.
-Therefore, the number of input and output channels of U-Net was set to 30 (15 coils).
-The proposed SGM-Net was trained on a UBUNTU 20.04 LTS system with a 3090 GPU (24GB) for 100 epoches, and the learning rate was set to 1e-3.
-The batch size is set to 1.
-V Experiments and Results
-V-A Dataset
-TABLE I: The quantitative results on Coronal PD and PDFS sequences, as well as the FastMRI dataset, compared with those of cutting-edge methods.
-Sequence Method PSNR SSIM
-Coronal PD zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
-TV 33.7498$\pm$3.1753 32.3290$\pm$3.2159 0.8851$\pm$0.0657 0.8658$\pm$0.0725
-U-Net 36.8984$\pm$2.7072 34.7352$\pm$2.7863 0.9361$\pm$0.0610 0.9133$\pm$0.0682
-D5C5 38.7903$\pm$2.8408 35.9697$\pm$2.7990 0.9509$\pm$0.0587 0.9259$\pm$0.0689
-ISTA-Net 39.2910$\pm$2.8423 35.5220$\pm$3.0915 0.9471$\pm$0.0567 0.9138$\pm$0.0717
-VS-Net 40.1649$\pm$2.8767 37.0604$\pm$2.7927 0.9591$\pm$0.0569 0.9358$\pm$0.0648
-E2E-VarNet 39.3592$\pm$2.9368 36.8949$\pm$3.3872 0.9542$\pm$0.0662 0.9342$\pm$0.0775
-ReVarNet 40.1141$\pm$2.8765 37.3681$\pm$2.7867 0.9590$\pm$0.0588 0.9379$\pm$0.0677
-MeDL-Net 40.9295$\pm$2.9072 38.1052$\pm$2.7622 0.9614$\pm$0.0557 0.9417$\pm$0.0663
-SGM-Net (ours) 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
-Coronal PDFS zero-filled 32.0299$\pm$2.0273 31.2221$\pm$2.0575 0.7964$\pm$0.0998 0.7587$\pm$0.1172
-TV 33.7256$\pm$1.9058 32.7328$\pm$1.8850 0.8124$\pm$0.0752 0.7920$\pm$0.0796
-U-Net 34.4367$\pm$0.6572 33.2028$\pm$2.5662 0.8183$\pm$0.0106 0.7758$\pm$0.1222
-D5C5 34.9671$\pm$2.8261 33.4189$\pm$2.4765 0.8250$\pm$0.1106 0.7854$\pm$0.1223
-ISTA-Net 34.4990$\pm$2.6389 33.4569$\pm$2.5758 0.8198$\pm$0.1066 0.7854$\pm$0.1214
-VS-Net 34.6847$\pm$2.7134 33.5010$\pm$2.6100 0.8259$\pm$0.1073 0.7842$\pm$0.1229
-E2E-VarNet 34.6274$\pm$2.7331 33.1626$\pm$2.4742 0.8234$\pm$0.1101 0.7804$\pm$0.1244
-ReVarNet 34.8522$\pm$2.7461 33.3146$\pm$2.5604 0.8272$\pm$0.1080 0.7832$\pm$0.1232
-MeDL-Net 35.5931$\pm$2.7307 34.1409$\pm$2.5501 0.8349$\pm$0.1025 0.7922$\pm$0.1255
-SGM-Net (ours) 35.7932$\pm$3.0718 34.3309$\pm$2.5192 0.8413$\pm$0.1076 0.7978$\pm$0.1242
-FastMRI zero-filled 33.2828$\pm$3.0594 32.5069$\pm$3.0657 0.8857$\pm$0.0709 0.8680$\pm$0.0791
-TV 35.5512$\pm$2.9189 33.8441$\pm$2.8703 0.8884$\pm$0.0635 0.0201$\pm$0.0152
-U-Net 35.21158$\pm$2.5925 33.7864$\pm$2.5331 0.8562$\pm$0.0574 0.8091$\pm$0.0637
-D5C5 36.4550$\pm$3.0246 34.5636$\pm$2.8510 0.9113$\pm$0.0649 0.8849$\pm$0.0749
-ISTA-Net 36.5753$\pm$2.9887 34.6229$\pm$2.8269 0.9134$\pm$0.0656 0.8837$\pm$0.0752
-VS-Net 37.3042$\pm$3.1130 35.5335$\pm$2.8513 0.9202$\pm$0.0637 0.8916$\pm$0.0736
-E2E-VarNet 37.5951$\pm$3.0129 35.6696$\pm$2.8326 0.9217$\pm$0.0635 0.8978$\pm$0.0726
-ReVarNet 37.0047$\pm$3.0130 34.8855$\pm$2.7926 0.9168$\pm$0.0646 0.8895$\pm$0.0746
-MeDL-Net 37.6791$\pm$2.9644 35.9594$\pm$2.7401 0.8984$\pm$0.0643 0.8882$\pm$0.0717
-SGM-Net (ours) 38.9212$\pm$3.4296 37.9430$\pm$3.2521 0.9344$\pm$0.0604 0.9234$\pm$0.0067
-In this study, we adopted a publicly available 15-coil knee dataset proposed in [51] to train our model and compared it with other methods.
-The knee dataset consisted of five 2D turbo spin-echo (TSE) sequences, each of which contained volumes scanned using a 3T scanner from 20 patients.
-We conducted experiments on coronal proton density (PD; TR=2750 ms, TE=27 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 35-42 slices), and coronal fat-saturated PD (PDFS; TR=2870 ms, TE=33 ms, in-plane resolution 0.49$\times$0.44 mm2, slice thickness 3 mm, 33-44 slices) sequences.
-The coil sensitivity maps were precomputed using ESPRIRiT and included in the dataset.
-The training, validation, and testing sets were randomly divided at a ratio of 3:1:1.
-Random Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration.
-8% lines at the center region are preserved.
-Examples of training images and the corresponding k-space trajectories are shown in Fig.
-3 (the first and second rows).
-FastMRI [8] is a publicly available dataset for accelerated MRI, which has been adopted in relevant studies.
-Considering the limitations of our hardware, we randomly selected a subset of the 15-coil knee dataset to evaluate our model and demonstrate its effectiveness.
-The data were acquired using the 2D TSE protocol, containing PD and PDFS sequence scanning from 3T or 1.5T systems (TR=2200-3000 ms, TE=27-34 ms, in-plane resolution 0.5$\times$0.5 mm2, slice thickness 3 mm).
-A total of 419, 33, and 375 slices were used for training, validation, and testing sets, respectively.
-Coil sensitivity maps were not available in the dataset, and related studies have been conducted to estimate and optimize sensitivity maps during training.
-However, we did not focus on SM estimation in this study and used pretrained SMs in our model to highlight the improvements brought about by the proposed methods.
-In the fastMRI dataset, we calculated the SMs using the sigpy [65] toolbox.
-Equispaced Cartesian Sampling masks are applied to obtain undersampled measurements with four-fold and six-fold acceleration.
-8% lines at the center region are preserved.
-Examples of training images and the corresponding k-space trajectories are shown in Fig.
-3 (the third and fourth rows).
-V-B comparisons with the cutting-edge methods
-We compared the proposed SGM-Net with two traditional and seven DL-based methods, as presented in Table I.
+We compared the proposed SGM-Net with two traditional and eight DL-based methods, as presented in Table 1.
 In the zero-fill method, the missing data in the measurements were filled with zero.
-The TV method uses total variation as the image prior in a CS-based approach.
+The TV method used total variation as the image prior in a CS-based approach.
 For DL-based methods (i.e., U-Net [8] [^2] https://github.com/facebookresearch/fastMRI, D5C5 [70] [^3] https://github.com/js3611/Deep-MRI-Reconstruction.
 ISTA-Net [52] [^4] https://github.com/jianzhangcs/ISTA-Net-PyTorch VSNet [54] [^5] https://github.com/j-duan/VS-Net.
-E2EVarNet [53] [^6] https://github.com/NKI-AI/direct, ReVarNet [67]6 and MeDL-Net [68] [^7] https://github.com/joexy312/MEDL-Net), we adopted the released codes and retrained the networks on our device.
+E2EVarNet [55] [^6] https://github.com/NKI-AI/direct, ReVarNet [71]6, MeDL-Net [72] [^7] https://github.com/joexy312/MEDL-Net and vSHARP [67]6), we adopted the released codes and retrained the networks on our device.
 We made a concerted effort to keep the announced settings same as those in the corresponding articles, except that we reduced the number of filters in ReVarNet so that all networks could be trained on one GPU.
 All methods were implemented on the knee dataset (coronal PD and PDFS sequences) and a subset of fastMRI dataset.
-The PSNR and SSIM results were calculated under four- and six-fold acceleration to evaluate the different methods.
-As shown in Table I, the proposed SGM-Net robustly achieves the best results on all the datasets with different acceleration factors.
-The reconstruction results of the sample images are presented in Fig.4.
+The PSNR and SSIM results were calculated under four- and six-fold acceleration to evaluate different methods.
+As shown in Table 1, the proposed SGM-Net robustly achieves the best results on all the selected datasets with different acceleration factors.
+Examples of reconstruction images are presented in Fig.4.
 We can observe from the zoomed region (second rows) that our method recovered the most details, and the error maps (third row) exhibited the least errors.
-V-C Ablation Studies
-TABLE II: Results of ablation studies on coronal PD sequence with four- and six-fold acceleration rates.
+5.3 Ablation Studies
+Table 2: Results of ablation studies on coronal PD sequence with four- and six-fold acceleration rates.
 Method PSNR SSIM
-SG 26.8198$\pm$4.7495 25.0439$\pm$7.2973 0.4172$\pm$0.1712 0.5212$\pm$0.1568
-MN 38.8362$\pm$2.6402 35.9488$\pm$2.6333 0.8623$\pm$0.0880 0.7861$\pm$0.1088
-SR-MN 35.4585$\pm$2.9562 33.2093$\pm$2.6398 0.8351$\pm$0.0912 0.7412$\pm$0.1284
-SG-MN-DM-US 40.7095$\pm$3.0165 38.0227$\pm$2.9647 0.9619$\pm$0.0615 0.9440$\pm$0.0700
-SG-MN-DM-DG 40.8443$\pm$3.0384 38.2183$\pm$2.9986 0.9629$\pm$0.0624 0.9448$\pm$0.0721
-SG-MN-US-DG 41.0424$\pm$3.4144 38.1616$\pm$2.9985 0.9633$\pm$0.0676 0.9443$\pm$0.0714
-SG-MN-DM-US-DG 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
-To highlight the improvements brought by different modules, we designed several workflows for a clear evaluation.
+SG 31.3478$\pm$2.6593 29.7293$\pm$5.1889 0.4730$\pm$0.1556 0.6067$\pm$0.1215
+MN 38.2080$\pm$2.4697 35.8126$\pm$2.4974 0.8346$\pm$0.1386 0.7849$\pm$0.1279
+SG-MN 35.4585$\pm$2.9562 33.2093$\pm$2.6398 0.8351$\pm$0.0912 0.7412$\pm$0.1284
+SG-MN-DM 39.3722$\pm$3.6295 36.9837$\pm$4.0466 0.9533$\pm$0.0801 0.9368$\pm$0.0812
+SG-MN-DM-US 40.5941$\pm$3.0464 37.8509$\pm$2.8824 0.9615$\pm$0.0619 0.9425$\pm$0.0684
+SG-MN-DM-DG 41.0969$\pm$3.7561 38.3457$\pm$2.9926 0.9625$\pm$0.0803 0.9440$\pm$0.0730
+SG-MN-US-DG 41.1213$\pm$3.3004 38.3419$\pm$3.0769 0.9630$\pm$0.0715 0.9444$\pm$0.0772
+SG-MN-DM-US-DG 41.1981$\pm$3.3802 38.4491$\pm$3.2166 0.9634$\pm$0.0682 0.9452$\pm$0.0790
+Table 3: Reconstruction results on coronal PD sequence with four- and six-fold acceleration rates using different inputs in DMs.
+structure Method PSNR SSIM
+w/o unrolled network $x_{T}$ 38.5337$\pm$3.8373 36.2220$\pm$4.5657 0.9487$\pm$0.0812 0.9322$\pm$0.0825
+$x_{T},x_{d}$ 38.9168$\pm$3.9265 36.4447$\pm$4.5595 0.9496$\pm$0.0808 0.9327$\pm$0.0818
+$x_{T},x_{p}$ 38.9123$\pm$3.4768 36.6218$\pm$4.1896 0.9507$\pm$0.0799 0.9345$\pm$0.0814
+$x_{T},x_{d},x_{p}$ 39.3722$\pm$3.6295 36.9837$\pm$4.0466 0.9533$\pm$0.0801 0.9368$\pm$0.0812
+w unrolled network $x_{T}$ 41.1164$\pm$3.4041 38.2711$\pm$3.6271 0.9629$\pm$0.0708 0.9441$\pm$0.0825
+$x_{T},x_{d}$ 41.0953$\pm$3.3920 38.2932$\pm$3.3948 0.9628$\pm$0.0669 0.9438$\pm$0.0814
+$x_{T},x_{p}$ 41.1620$\pm$3.1978 38.4075$\pm$3.3948 0.9633$\pm$0.0639 0.9450$\pm$0.0812
+$x_{T},x_{d},x_{p}$ 41.1981$\pm$3.3802 38.4491$\pm$3.2166 0.9634$\pm$0.0682 0.9452$\pm$0.0790
+To highlight the improvements brought by different modules, we designed several ablation workflows for a clear evaluation.
 Specifically, the proposed modules for evaluation were the SMLD-based guidance (SG), model-driven network structure (MN), denoising module (DM), updating steps (US) for SGs, and densely connected guidance (DG) in the unrolled network.
@@
   2. 2.
-    MN: a standard unrolled network with the undersampled measurements as the inputs.
+    MN: a standard unrolled network without guidances. The undersampled measurements are adopted as the inputs.
   3. 3.
@@
   4. 4.
+    SG-MN-DM: The DGIs in SGM-Net are adopted as the final reconstructions.
+  5. 5.
     SG-MN-DM-US: the densely connected structure is removed from all the cascades.
-  5. 5.
-    SG-MN-DM-DG: the SMLD results are used as guidance in the unrolled network but without and updating steps.
   6. 6.
-    SG-MN-US-DG: the SMLD results are used to guide the network without denoising in the DM.
+    SG-MN-DM-DG: the SMLD results are used as guidance in the unrolled network but without the updating steps.
   7. 7.
+    SG-MN-US-DG: the SMLD results are used to guide the network without denoising by the DM.
+  8. 8.
     SG-MN-DM-US-DG: the proposed SGM-Net.
-The quantitative results are shown in table II.
-Comparing SG with the proposed SGM-Net, we can observe a significant improvement brought about by the proposed modules, whereas the average results from the SMLD sampling are worse than those of the zero-filled method.
-Similarly, through the guidance of the SMLD results, the results of the unrolled MN method were significantly improved with guidance from the SMLD results.
-Using either of these two methods alone can’t achieve satisfactory results.
-Moreover, we simply combine SR and MN in the SR-MN workflow that the SMLD results are directly fed to the network as input.
-However, the performance are even worse than that of MN.
-This phenomenon demonstrates the effectiveness of the proposed guidance-based workflow, which outperforms a simple combination of the two methods.
-A comparison between SG-MN-DM-US and SGM-Net demonstrated that the densely connected structure is capable of fulfilling the guiding work and improving the final results.
-Removing the densely connected structure and using only the AMs to optimize the reconstructed images will lead to a degredation in the final results.
-By comparing SG-MN-DM-DG with SGM-Net, we observed that the updating steps in the proposed model can ensured that accurate guidance was provided by the PGIs and improved the results.
-When removing the DM and directly feeding SMLD results to guide the network, the reconstruction performances of SG-MN-US-DG are similarly worse than that of SGM-Net.
-The best results are achieved with the integration of all the proposed modules, which demonstrates their effectiveness and necessity.
-Furthermore, we illustrate in subsection V.E that the integration of the proposed module can enhance the robustness of the network with greatly reduced training data.
-Figure 5: The reconstruction results of the original model at four-fold acceleration.
-The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps representing the objective reality.
-The corresponding images from left to right: the objective reality, zero-filled measurements, PGI, DGI, and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branch.
-Figure 6: The reconstruction results with 20% sampling steps at four-fold acceleration.
-The first rows of each subfigure are the reconstructed images, the second rows are zoomed details in the red square, and the third rows are the error maps representing the objective reality.
-The corresponding images from left to right: the objective reality, zero-filled measurements, PGI, DGI, and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branch.
-TABLE III: Quantitative results of the selected images at four-fold and six-fold acceleration rates.
+The quantitative results are shown in table 2.
+Comparing SG with the proposed SGM-Net, we observed a significant improvement brought by the proposed modules, whereas the average results from the SMLD sampling were even worse than those of the zero-filled method.
+Similarly, through the guidance of the SMLD results, the results of the unrolled MN method were significantly improved.
+Using either of these two methods alone could not achieve satisfactory results.
+Moreover, we simply combined SG and MN in the SG-MN workflow that the SMLD results were directly fed to the network as input.
+However, the performance was worse than that of MN.
+This phenomenon demonstrated the effectiveness of the proposed guidance-based workflow, which outperformed a simple combination of the two methods.
+Comparing with the proposed SGM-Net, the SG-MN-DM workflow removed the guided reconstruction structures, and we observed deterioration on the results.
+On the other hand, the improvements brought by the DMs over SG and MN method were also obvious.
+A comparison between SG-MN-DM-US and SGM-Net demonstrated that the densely connected structure was capable of fulfilling the guiding work and significantly improving the final results.
+Removing the densely connected structure and using only the AMs to optimize the reconstructed images would lead to a degredation in the final results.
+By comparing SG-MN-DM-DG with SGM-Net, we observed that the updating steps in the proposed model can further ensured that accurate guidance was provided by the PGIs and improved the results.
+When removing the DM and directly feeding SMLD results to guide the network, the reconstruction performances of SG-MN-US-DG were similarly worse than that of SGM-Net.
+Meanwhile, when removing DM or US structure from SGM-Net, we observed that the network still achieved convincing results.
+Specifically, when removing US, the denoised output of DM can provide effective guidance for the subsequent reconstructions.
+When removing DM, the updating steps in the networks improved the image quality of SMLD results during training.
+This phenomenon further demonstrated the robustness of the proposed network structure.
+The best results were achieved with the integration of all the proposed modules, which demonstrated their effectiveness and necessity.
+Furthermore, we illustrate in subsection 5.5 that the integration of the proposed module can enhance the robustness of the network with greatly reduced training data.
+Besides, we conducted experiments using different features in DMs, and the results are shown in Table 3.
+Here, we evaluated the results with and without the guided reconstruction process.
+When the unrolled network structure was removed, the improvements brought by involving $x_{d}$ or $x_{p}$ in DMs is evident, which clearly exhibited the improvements brought by the SIE and CIE blocks.
+With the unrolled network structure, the distances among the methods were narrowed, that the unrolled network can compensate for the deterioration in DMs.
+However, incorporating all the components in DMs exhibited the best results, demonstrating the necessity of the proposed SIE and CIE blocks.
+Figure 5: (a) The reconstruction results of the original model at four-fold acceleration.
+(b) The reconstruction results with 20% sampling steps at four-fold acceleration.
+The first rows of each subfigure are the reconstructed images, and the second rows are zoomed details in the red squares.
+The corresponding images from left to right: the ground truth, zero-filled measurements, PGI, DGI, and the final output of $x_{T}^{K}$ and $x_{z}^{K}$ branch.
+Figure 6: (a) The reconstruction results on fastMRI dataset at four-fold acceleration.
+(b) The reconstruction results on fastMRI dataset at six-fold acceleration.
+The first rows of each subfigure are the reconstructed images and the second rows are zoomed details in the red squares.
+The corresponding images from left to right: the ground truth, zero-filled measurements, PGI, DGI, and the final output of $x_{T}^{K}$ and $x_{z}^{K}$ branch.
+Table 4: Quantitative results of the selected images at four-fold and six-fold acceleration rates.
 Method Image PSNR SSIM
-zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
-original $x_{T}$ 26.8198$\pm$4.7495 25.0439$\pm$7.2973 0.4172$\pm$0.1712 0.5212$\pm$0.1568
-$x_{T}^{0}$ 39.0871$\pm$2.5214 37.0431$\pm$2.3824 0.9523$\pm$0.0587 0.9364$\pm$0.0635
-$x_{T}^{K}$ 41.0913$\pm$3.4621 38.3504$\pm$2.9553 0.9635$\pm$0.0656 0.9458$\pm$0.0678
-$x_{z}^{K}$ 41.2213$\pm$3.3385 38.4681$\pm$2.9889 0.9638$\pm$0.0642 0.9459$\pm$0.0678
-20% sampling steps $x_{T}$ 19.7481$\pm$7.8121 17.5994$\pm$7.7102 0.4664$\pm$0.2049 0.4312$\pm$0.1723
-$x_{T}^{0}$ 38.4951$\pm$2.9590 36.4403$\pm$2.9405 0.9500$\pm$0.0600 0.9317$\pm$0.0656
-$x_{T}^{K}$ 40.8341$\pm$3.8591 38.2336$\pm$3.0403 0.9631$\pm$0.0654 0.9421$\pm$0.0691
-$x_{z}^{K}$ 40.9907$\pm$3.7419 38.3414$\pm$2.8912 0.9631$\pm$0.0657 0.9446$\pm$0.0682
-pretrained with another dataset $x_{T}$ 25.5783$\pm$4.7741 24.2301$\pm$7.8034 0.4904$\pm$0.1506 0.4692$\pm$0.1549
-$x_{T}^{0}$ 39.0112$\pm$2.7413 36.8603$\pm$2.7916 0.9510$\pm$0.0578 0.9351$\pm$0.6225
-$x_{T}^{K}$ 40.9581$\pm$3.3235 38.2714$\pm$2.8405 0.9634$\pm$0.0663 0.9401$\pm$0.0702
-$x_{z}^{K}$ 41.0920$\pm$3.2329 38.4204$\pm$2.9084 0.9636$\pm$0.6828 0.9432$\pm$0.0688
-V-D Evaluation of the guided reconstruction process
-To evaluate the improvements brought by the guided process, we plotted the following images: the image of the ground truth, zero-filled measurements, PGI ($x_{T}$), DGI ($x_{T}^{0}$), and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branches ($x_{T}^{K}$ and $x_{z}^{K}$) of an MR image from the test dataset.
-To mitigate potential influences from the sampling process, we designed the following three sampling methods:
-  1. 1.
-    original: the SMLD results are sampled using the pretrained score network, as described previously.
-  2. 2.
-    20% sampling steps: the SMLD results are sampled using the pretrained score network, but using only 20% of the original sampling steps.
-  3. 3.
-    pretrained with another dataset: we continued to train the score network on another brain dataset proposed in [69], and the images are sampled with the same hyperparameters in the original method.
+original, Coronal PD zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
+$x_{T}$ 31.3478$\pm$2.6593 29.7293$\pm$5.1889 0.4730$\pm$0.1556 0.6067$\pm$0.1215
+$x_{T}^{0}$ 38.9772$\pm$3.6659 37.0103$\pm$4.1097 0.9508$\pm$0.0798 0.9370$\pm$0.0812
+$x_{T}^{K}$ 41.1457$\pm$3.6320 38.2893$\pm$3.9790 0.9630$\pm$0.0750 0.9450$\pm$0.0828
+$x_{z}^{K}$ 41.1981$\pm$3.3802 38.4491$\pm$3.2166 0.9634$\pm$0.0682 0.9452$\pm$0.0790
+20% sampling steps, Coronal PD zero-filled 31.3089$\pm$3.3939 30.5176$\pm$3.3902 0.8778$\pm$0.0713 0.8626$\pm$0.0799
+$x_{T}$ 22.5488$\pm$7.6208 20.6821$\pm$7.8592 0.4664$\pm$0.2049 0.4601$\pm$0.2810
+$x_{T}^{0}$ 37.8958$\pm$4.2813 36.6910$\pm$4.1992 0.9461$\pm$0.0809 0.9326$\pm$0.0837
+$x_{T}^{K}$ 40.9129$\pm$4.2478 38.2178$\pm$3.8602 0.9622$\pm$0.0817 0.9403$\pm$0.0826
+$x_{z}^{K}$ 40.9847$\pm$3.6107 38.2610$\pm$3.6829 0.9628$\pm$0.0770 0.9412$\pm$0.0798
+original, FastMRI zero-filled 33.2828$\pm$3.0594 32.5069$\pm$3.0657 0.8857$\pm$0.0709 0.8680$\pm$0.0791
+$x_{T}$ 26.5182$\pm$5.1008 23.9689$\pm$5.0221 0.4081$\pm$0.1377 0.4533$\pm$0.1307
+$x_{T}^{0}$ 37.3427$\pm$3.5911 35.0240$\pm$3.3366 0.9236$\pm$0.0682 0.8933$\pm$0.0799
+$x_{T}^{K}$ 38.7848$\pm$3.4290 36.8528$\pm$3.0269 0.9338$\pm$0.0591 0.9092$\pm$0.0705
+$x_{z}^{K}$ 38.8241$\pm$3.3808 36.8730$\pm$3.0364 0.9344$\pm$0.0588 0.9094$\pm$0.0707
+5.4 Evaluating the effectiveness of the proposed workflow
+To evaluate the effectiveness of the proposed workflow, we plotted the following images: the ground truth, zero-filled measurements, PGI ($x_{T}$), DGI ($x_{T}^{0}$), and the final output of $x_{z}^{k}$ and $x_{T}^{k}$ branches ($x_{T}^{K}$ and $x_{z}^{K}$) of an MR image from the test dataset.
 The reconstructed images from the original sampling method at four-fold acceleration are shown in Fig.
-5.
+5 (a).
 The first row shows the corresponding images, and the second row shows the magnified details.
-The zero-filled image (second column) is blurred, and the PGI (third column) loses texture details.
-However, the artifacts in the DGI (fourth column) were effectively eliminated from the zero-filled image and PGI by DMs.
-Then, the GICs recovered the missing details in the DGIs and $x_{T}^{K}$ (fifth column) and $x_{z}^{K}$ (sixth column) presented satisfactory results when compared with the ground truth images (first column).
-The error maps in the third row exhibit that the massive errors in the zero-filled images and PGIs are obviously alleviated, the the image quality are further improved in $x_{T}^{K}$ and $x_{z}^{K}$.
+We can observe that the zero-filled image (second column) is blurred, and the PGI (third column) exhibits hallucination artifacts.
+However, the artifacts in the DGI (fourth column) were effectively eliminated by DMs from the zero-filled image and PGI.
+Then, the GICs further enhanced the image quality of the DGIs, and $x_{T}^{K}$ (fifth column) and $x_{z}^{K}$ (sixth column) presented satisfactory results when compared with the ground truth images (first column), demonstrating the proposed coarse-to-fine workflow has powerful ability in recovering fine details.
+Besides, we sampled PGIs from the pretrained score network, but using only 20% of the original sampling steps.
 The reconstructed images with 20% sampling steps at four-fold acceleration are depicted in Fig.
-6, which are also consistent with the observations in Fig.
-5.
-The quantitative results for the three sampling methods are presented in Table III.
+5 (b).
+Adopting PGIs with worse quality, the proposed method still robustly achieve high-quality results, which are also consistent with the observations in Fig.
+5 (a).
+Besides, we plotted reconstructed images from the FastMRI dataset at four-fold and six-fold in Fig.
+6 (a) and (b).
+We observed unexpected artifacts on the PGIs, which lower the image quality.
+However, through the denoising steps, we obseved that the artifacts have been greatly suppressed in DGIs, and the outputs of the unrolled network further eliminated the influence of noise and artifacts.
+The quantitative results for the abovementioned methods are presented in Table 4.
 $x_{z}^{K}$, which is chosen as the final result, slightly outperforms $x_{T}^{K}$, and the other results obtained at four- or six-fold acceleration are consistent with the results observed from the images.
@@
 However, the impact on the final reconstruction results is minimal, demonstrating the robustness of our proposed method to noise input, and the DMs are effective to extract valuable information from the noise input.
-We continue training the score network on another brain dataset to make perturbations on the pretrained model, which produces reconstructions of equally high quality through our workflow, demonstrating that our method imposes minimal requirements on the sampling process.
-In summary, this experiment vividly illustrates how our proposed workflow leverages SMLD results to extract crucial information, aiding in the reconstruction process and finally achieving high-quality results.
-V-E Evaluation of the network robustness
-Figure 7: PSNR (the first column) and SSIM (the second column) results of SGM-Net, MeDL-Net, and E2EVarNet with different amounts of training data at four- (the first row) and six-fold (the second row).
-In this subsection we illustrate the network robustness with limited training data.
-Here, we compared E2E-VarNet and MeDL-Net, which have convincing performances and do not leverage guidance images, as shown in Table I.
+In summary, this experiment vividly illustrates how the coarse-to-fine workflow leverages SMLD results to extract crucial information, aiding in the reconstruction process and finally achieving high-quality results.
+5.5 Evaluating the network robustness with limited training slices
+Figure 7: Four- and six-fold PSNR and SSIM results of different networks with limited training data.
+In this subsection we evaluated the network robustness with limited training data.
+Here, we first compared with E2E-VarNet and MeDL-Net, which have shown convincing performances without utilizing guidance images.
+We compared them with SGM-Net on the coronal PD sequence, and the PSNR and SSIM results at four-fold and six-fold are depicted in Fig 5.5 (a)-(d).
+We trained the networks with 20, 50 and 100 slices, which are randomly selected, to evaluate their robustness.
+We observed that the performances of E2E-VarNet and MeDL-Net deteriorated with decreasing training data.
+Meanwhile, the performance of SGM-Net was more stable with greatly reduced training slices.
 We also compared with three ablation networks (i.e.
-SG-MN-DM-US, SG-MN-MN-DG and SG-MN-US-DG).
-We reduced the number of training images of the different methods to different extents and compared their performances with our model.
-The results are presented in Fig.
-7.
-We compared the PSNR and SSIM results of the above networks at four- and six-fold, respectively.
-We observed that the performances of the other networks deteriorated with decreasing training data.
-Meanwhile, the performance of SGM-Net was more stable even with only 20 training images.
-Compared with E2E-VarNet and MeDL-Net, our ablation networks also exhibit relatively better robustness.
-However, the degradation of image quality become more pronounced with the decrease of training data, which further demonstrating the necessity of the proposed modules.
-To sum up, this experiment demonstrates that proposed guidance process can effectively compensate for information in the missing training data and make the model robust.
-VI Discussions
+SG-MN-DM-US, SG-MN-MN-DG and SG-MN-US-DG), and the results are depicted in Fig 5.5 (e)-(h).
+Compared with E2E-VarNet and MeDL-Net, the ablation networks similarly showed better robustness, with a slower degradation in image quality under limited training data conditions, which also indicates the robustness of the proposed workflow.
+Furthermore, the degradation of image quality become more pronounced with the decrease of training data, which demonstrated the necessity of the proposed modules.
+To sum up, this experiment demonstrated that proposed guided reconstruction process can effectively compensate for information in the missing training data and achieve convincing performance robustly.
+6 Discussions
 To make the sampling process robust and achieve better performance, most of the cutting-edge methods proposed to optimize the training of the generative models or the sampling steps.
@@
 During inference, the reconstruction process iterates between SDE solvers and data consistency steps.
-Compared with them, the most significant distinction in our workflow is that we bypass all the tedious hyperparameter tuning and network retraining steps and implement a naive sampling.
+In [15], the authors enhanced the reconstruction robustness of diffusion models with test-time hyperparameter tuning based on Stein’s unbiased risk estimator.
+Compared with them, the most significant distinction in our workflow is that we bypassed all the tedious hyperparameter tuning and network retraining steps and implemented a naive sampling.
 The reconstruction performances are then not robust especially with distribution-shift test datasets, and the results are observed to be severely corrupted.
-We novelly treat the SMLD results as a generalized prior, and we design an unrolled network which can extract valuable information from the SMLD samples to guide reconstruction, while the components during sampling can also help to eliminate the artifacts in the network.
+We treat the SMLD results as a generalized prior, and we designed an unrolled network which can extract valuable information from the SMLD samples to guide reconstruction.
 Experimental results have demonstrated that compared with the naive sampling results, only small amounts of data are required to train the unrolled network and the performances can be greatly improved.
 Another bottleneck in diffusion models is that the sampling process is relatively slow.
-To solve this problem, the authors in [70] proposed DiffuseRecon, and the reverse process can be guided with observed k-space signal.
+To solve this problem, the authors in [73], for example, proposed DiffuseRecon, and the reverse process can be guided with observed k-space signal.
 The authors designed a coarse-to-fine algorithm to save sampling time.
@@
 However, the unrolled network have good tolerance to artifact-affected inputs, hence the sampling steps can also be great reduced.
-In the proposed reconstruction workflow, the low-quality SMLD samples have been demonstrated to be effective to guide the network training and improve the final results.
+In the proposed reconstruction workflow, the coarse SMLD samples have been demonstrated to be effective to guide the network training and improve the final results.
 Meanwhile, the proposed workflow is also compatible to optimized SMLD steps with higher quality references, which show great potential to be exploited in future works.
@@
 Training strategies such as self-supervising should be considered to further optimize the workflow in future works.
-VII Conclusions
-In this study, we proposed a novel workflow to eliminate the need for optimizing the forward and reverse process in the SMLD-based method for MRI reconstruction.
-Instead, an unrolled model-driven network is proposed to extract valuable information from naive sampling results.
-Subsequently, the SMLD results can be periodically updated and utilized to guide the network training process.
-The experimental results demonstrated that the proposed hybrid method outperforms the workflows solely based on diffusion models or unrolled networks.
+7 Conclusions
+In this study, we proposed a novel MRI reconstruction workflow consisting of sampling, denoising and guided reconstruction steps.
+Naive SMLD samples are denoised in DMs and the artifacts are effectively mitigated.
+Then, the denoised samples are periodically updated and utilized to guide the network training process to recover fine details.
+Experimental results demonstrated that the proposed method outperforms the workflows solely based on diffusion models or unrolled networks.
 Additionally, our method alleviates the substantial requirements for sampling steps and training images.
 When compared with cutting-edge methods, our network consistently produces reconstructions of higher quality, showcasing its superiority in MRI reconstruction tasks.
+8 Acknowledgements
+This work was supported by the National Natural Science Foundation of China [Nos.
+62331008, 62027827, 62221005 and 62276040], National Science Foundation of Chongqing [Nos.
+2023NSCQ-LZX0045 and CSTB2022NSCQ-MSX0436], Chongqing University of Posts and Telecommunications ph.D.
+Innovative Talents Project [BYJS202211].